### v1

# Breaking the Curse of Dimensionality with Distributed Neural Computation

Abstract
We present a theoretical approach to overcome the curse of dimensionality using a neural computation algorithm which can be distributed across several machines. Our modular distributed deep learning paradigm, termed neural pathways, can achieve arbitrary accuracy while only loading a small number of parameters into GPU VRAM. Formally, we prove that for every error level $\varepsilon>0$ and every Lipschitz function $f:[0,1]^{n}\to\mathbb{R}$, one can construct a neural pathways model which uniformly approximates $f$ to $\varepsilon$ accuracy over $[0,1]^{n}$ while only requiring networks of $\mathcal{O}(\varepsilon^{-1})$ parameters to be loaded in memory and $\mathcal{O}(\varepsilon^{-1}\log(\varepsilon^{-1}))$ to be loaded during the forward pass. This improves the optimal bounds for traditional non-distributed deep learning models, namely ReLU MLPs, which need $\mathcal{O}(\varepsilon^{-n/2})$ parameters to achieve the same accuracy. The only other available deep learning model that breaks the curse of dimensionality is MLPs with super-expressive activation functions. However, we demonstrate that these models have an infinite VC dimension, even with bounded depth and width restrictions, unlike the neural pathways model. This implies that only the latter generalizes. Our analysis is validated experimentally in both regression and classification tasks, demonstrating that our model exhibits superior performance compared to larger centralized benchmarks.
Distributed Computing, Approximation Theory, Deep Learning, VC Dimension.
1 Introduction
The curse of dimensionality is a core obstruction in the foundations of machine learning. From the theoretical lens, given by the universal approximation theorem, the curse of dimensionality implies that every standard deep learning model approximately implementing an arbitrary almost-everywhere differentiable function on the $n$-dimensional cube $[0,1]^{n}$, up to an error of $\varepsilon>0$, requires at least $\Omega(\varepsilon^{-n/2})$ parameters (Yarotsky, 2018; Kratsios & Papon, 2022; Shen et al., 2021, 2022b). In practice, model upscaling is limited by the available device memory; Video RAM (VRAM) for GPUs. This happens either in training, during the backward pass when computing the Jacobian of the deep learning model, or even in the forward pass if the model is large enough. Currently, the curse of dimensionality is only known to be breakable in either one of two restrictive scenarios:
Table 1: Overview of Main Results: Depth and width of the MLP architecture $\hat{f}_{\theta}$ and the number of active parameters in GPU VRAM (Params. in GPU) required to approximate an $\alpha$-Hölder function. The maximum number of parameters required during a forward pass, for any given input, are also recorded (Forward Pass). The optimal approximation rates for the standard ReLU MLP are derived in Yarotsky (2017); the optimality of which is expressed in terms of VC dimension in Shen et al. (2022b). The rates for the non-Lipschitz continuous “unstable MLP model” are derived in Shen et al. (2022a). The VC dimension of the class of ReLU MLPs with bounded depth and width is computed in Bartlett et al. (2019), the (finite) VC dimension of our neural pathways model, with bounded depth, width, and number of parts, is shown in Proposition 3.6, and the infinite VC dimension of the class of unstable MLPs is shown in Proposition 3.7.
Model Depth Width Params. in (V)RAM Forward Pass VC Dim Ours $\mathcal{O}(\varepsilon^{-1})$ $\mathcal{O}(1)$ ${\color[rgb]{0.09,0.45,0.27}\definecolor[named]{pgfstrokecolor}{rgb}{0.09,0.45,0.27}\pgfsys@color@rgb@stroke{0.09}{0.45}{0.27}\pgfsys@color@rgb@fill{0.09}{0.45}{0.27}{\mathcal{O}(\varepsilon^{-1})}}$ $\mathcal{O}\big{(}\frac{n}{\alpha}\log(1/\varepsilon)/\varepsilon\big{)}$ Finite ReLU MLP ${\color[rgb]{0.82,0.1,0.26}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.1,0.26}\pgfsys@color@rgb@stroke{0.82}{0.1}{0.26}\pgfsys@color@rgb@fill{0.82}{0.1}{0.26}{\mathcal{O}(\varepsilon^{-n/(2\alpha)})}}$ $\mathcal{O}(1)$ ${\color[rgb]{0.82,0.1,0.26}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.1,0.26}\pgfsys@color@rgb@stroke{0.82}{0.1}{0.26}\pgfsys@color@rgb@fill{0.82}{0.1}{0.26}{\mathcal{O}\big{(}\varepsilon^{-\frac{n}{2\alpha}}\big{)}}}$ $\mathcal{O}\big{(}\varepsilon^{-\frac{n}{2\alpha}}\big{)}$ Finite Unstable MLP $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Infinite
Exceptionally Regular Target Functions. The curse of dimensionality can only be broken when approximating members of highly restrictive classes of target functions (Barron, 1993; Mhaskar et al., 2017; Suzuki, 2019; Opschoor et al., 2022; Galimberti et al., 2022; Gonon et al., 2023). However, there are generally no verifiable guarantees that a target function encountered in practice has the necessary structure for these desired approximation theorems to hold.
Generalization Problems for Super-expressive Multilayer Perceptions (MLPs). Deep learning models with highly oscillatory “super-expressive” activation functions (Yarotsky & Zhevnerchuk, 2020; Yarotsky, 2021; Zhang et al., 2022) boast approximation rates which are not cursed by dimensionality. However, as we will see in Proposition 3.7, many of these models have an infinite VC dimension even when they are restricted to having a bounded depth and width. Their unbounded VC dimension implies that the classifiers implemented by these models, by post-composition with $I_{(0,\infty)}$, do not generalize due to the classical results of Blumer et al. (1989); Hanneke (2016); Brukhim et al. (2022). Thus, the real performance of these models need not achieve the approximation-theoretic optima since they can learn from a finite number of training instances.
This paper demonstrates that distributing neural computation offers a feasible third option for overcoming the curse of dimensionality in deep learning. Importantly, unlike the aforementioned cases, we show that distributed deep learning does not require high regularity of the target function being approximated, nor does it rely on models with unorthodox activation functions. Our deep learning algorithm, termed neural pathways, is distributed across the storage space (i.e., offline memory) of the machine. It only requires a small portion of the model to be loaded into VRAM at any given time, while the entire model is distributed over the machine’s (pragmatically infinite) storage space.
In essence, each point in the input space is assigned to a trainable prototype (or landmark or representative point). Each of these prototypes is linked with a small neural network, specifically designed to approximate the target function for inputs in proximity to its associated prototype. During the forward pass, the input is initially directed to the appropriate prototype, after which the corresponding network is loaded into memory and employed to generate a prediction. Our algorithm follows the principle of route-first and predict-second via locally specialized neural networks. This approach mimics how stimuli activate only specialized neural pathways in the brain, without necessitating the involvement of the entire brain to process any single input.
2 Neural Pathways
In this section, we present the motivation behind the neural pathways framework proposed in this paper, and we describe it from an algorithmic perspective. The primary goal is to provide the reader with an intuitive understanding of the overall procedure before delving into a more involved mathematical analysis in Section 3, which is approached through the lens of approximation theory.
Motivation. Our model formulation is loosely inspired by the cognitive sciences, and more specifically, by the fact that neuroscientists have observed that the brain does not activate completely to process each input, but rather uses specific regions of the overall network at a time, also known as brain functional specialization (Mahon & Cantlon, 2011).
On the other hand, from a computational perspective, our motivation is to increase the size of deep learning models regardless of GPU memory capacity. We theoretically study the effect and feasibility of scaling up models by combining smaller distributed networks specialized to approximate specific regions of the input space. This differs from simply increasing the parameter count of a single large network, which is constrained by available VRAM.
2.1 Model overview
We can visualize our framework as a tree, with each leaf corresponding to a prototype and an associated neural network. In our case, we work with MLPs, but the network need not be a specific architecture. We refer to the combination of each prototype and its associated neural network as a neural pathway within our model. We refer to the overall algorithmic construction as neural pathways in plural. We can think of neural pathways as a framework for neural computation rather than a specific model since this approach could be applied to any other common neural networks such as CNNs (LeCun et al., 1989) or Transformers (Vaswani et al., 2017), for instance. Similarly, we propose specific algorithms to discover the prototypes for our neural pathways, but these could be substituted with enhanced approaches in the future. However, the main goal of this work is not to adapt the framework to a variety of specialized machine learning tasks but rather to provide its initial formulation and, particularly, to analyze its mathematical feasibility in terms of the curse of dimensionality.
The neural pathways algorithm is divided into two stages: first, we discover the prototype locations, and then, we optimize the neural pathways associated with each prototype. In step one, our framework learns how to subdivide the input space, and in step two, we focus on neural network parameter optimization while keeping the input space partitioning fixed. The mathematical soundness of this approach is later presented as part of our main theoretical results in Section 3. A schematic of the algorithm workflow is depicted in Figure 1.
Figure 1: $1$) The distance from each input $x$ to all prototypes $p_{1},\dots,p_{8}$ is queried. $2$) The network ($\hat{f}_{2}$ in the figure) assigned to the nearest prototype ($p_{2}$ in the figure) is loaded onto the GPU and used for prediction.
2.2 Training Algorithm
We now provide an explanation for the training algorithm. As discussed, we break down the algorithm into two parts: discovering the prototypes and training the networks. Conceptually, the prototypes define where in the input space the networks are located, or in other words, where in the input space we expect each of the networks to have the best performance. During inference, we will route a given input to the appropriate network based on its nearest prototype. In essence, each network learns to approximate a specific region of the overall input domain.
Discovering prototypes. In principle, we may not know how to partition the input space. One approach is to utilize standard clustering algorithms like $K$-means, but this might be suboptimal for the downstream task unless we are already operating in a structured latent space, such as those found in pre-trained models (further discussion on this is available in Section 5). Another way is to learn it via gradient descent by optimizing the location of the prototypes for a specific task by following the gradient of the downstream loss. At the beginning of training, we have $\hat{F}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\hat{f}_{1}(x),\dots,\hat{f}_{K}(x))$ which contains a collection of $K$ randomly initialized shallow or small networks (i.e., much smaller than our neural pathways described later). In this first step, we assume that we are able to load all randomly initialized networks into our GPU memory. In particular, this is true because we use small networks with few parameters, which we will later “deepen” in the next step by adding additional hidden layers. We initialize the prototypes $p\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(p_{1},\dots,p_{K})$ randomly from a uniform distribution within the bounds of our training dataset input samples. We use the following expression to train the location of our prototypes $\{p_{k}\}_{k=1}^{K}$ in $\mathbb{R}^{n}$ by minimizing the energy:
$\sum_{(x,y)\in\mathcal{D}}\,\ell\big{(}\operatorname{softmax}\big{(}-\|x-p_{i}\|_{i=1}^{K}\big{)}^{\top}\,\hat{F}(x)\,,y\big{)}.$ (1)
where the loss $\ell$ is task-specific; for example, one could use mean squared error for regression and cross-entropy for classification. The $\operatorname{softmax}$ weights the importance of the prediction of each of the neural pathways in $\hat{F}$ for a given input $x$, as a function of the input’s distance to the prototypes, $\|x-p_{i}\|_{i=1}^{K}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\|x-p_{1}\|,...,\|x-p_{K}\|)$. Both the locations of the prototypes and the shallow randomly initialized neural networks assigned to them are optimized.
Deepening the Networks. After the initial training phase, we enhance the networks by incorporating additional layers. Specifically, we introduce linear layers with weights initialized to the identity matrix and bias set to zero, just before the final output layer of each network. To encourage gradient flow in these new layers during the subsequent training stage, we slightly perturb this initialization with small Gaussian noise. This approach is driven by the fact that in the second training stage, each neural pathway can be optimized in a distributed manner. Consequently, we can work with larger networks without the need to load all of them simultaneously into our GPU, allowing for more model parameters. During the first stage, we have already optimized our networks alongside the prototype locations, converging towards a minimum. By initializing the networks with the new layers close to the identity, we can ensure that their output at the start of the second stage of training is similar to that produced by the original networks. This allows us to smoothly continue the optimization process from the point where we previously halted.
Neural Pathway Training. Once prototype locations have been fixed we can independently train neural pathways $\hat{f}_{1},\dots,\hat{f}_{K}$ by minimizing for all $k\in\{1,\dots,K\}$:
($\underset{\underset{k\in\operatorname*{arg\,min}_{j\in\{1,\dots,K\}}\{\|x-p_{j}\|\}}{(x,y)\in\mathcal{D}}}{\sum}\,\ell(\hat{f}_{k}(x),y)$ (2)
over all the networks. We optimize the neural pathway network $\hat{f}_{k}$ for training data points that are closest to prototype $p_{k}$. The training procedure is summarized in Algorithm 1.
Algorithm 1 Neural Pathways Training.
Training data $\mathcal{D}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(x_{j},y_{j})\}_{j=1}^{N}$, no. of prototypes $K\in\mathbb{N}_{+}$, loss function $\ell$.
Discovering Prototypes:
$(\hat{F},p)\leftarrow{\operatorname*{arg\,min}_{\hat{F},p}}\,\displaystyle\sum_{(x,y)\in\mathcal{D}}\,\ell\left(\operatorname{softmax}\left(x|p\right)^{\top}\hat{F}(x),y\right)$
Deepen networks:
    For $k=1,\dots,K$:
     $\hat{f}_{k}\leftarrow\textrm{deepen}(\hat{f}_{k})$
Neural Pathway Training:
   For $k=1,\dots,K$:
     $\hat{f}_{k}\leftarrow\underset{\hat{f}_{k}}{\operatorname*{arg\,min}}\,\underset{\underset{k\in\operatorname*{arg\,min}_{j}\{\|x-p_{j}\|\}}{(x,y)\in\mathcal{D}}}{\sum}\,\ell(\hat{f}_{k}(x),y)$
return Neural pathway parameters $\{\hat{f}_{k}\}_{k=1}^{K}$ and prototype locations $\{p_{k}\}_{k=1}^{K}$.
Inference. At inference time, each test sample $x$ is assigned to its nearest prototype $p_{k}$ where $k\in\operatorname*{arg\,min}_{j\in\{1,\dots,K\}}\{\|x-p_{j}\|\}$ and the prediction is made by the $k$-th neural pathway $\hat{f}_{k}$.
3 Mathematical Analysis of Neural Pathways
We now analyze the approximation capacity of the neural pathways framework, when applied to MLPs with trainable PReLU activation functions.
3.1 Preliminaries
Before presenting our main result, Theorem 3.4, we first formalize the necessary mathematical terminology.
Notation. We use the following notation: for any $f,g:\mathbb{R}\to\mathbb{R}$, we write $f\in\mathcal{O}(g)$ if there exist $x_{0}\in\mathbb{R}$ and $M\geq 0$ such that for each $x\geq x_{0}$ we have $|f(x)|\leq Mg(x_{0})$. Similarly, we write $f\in\Omega(g)$ to denote the relation $g\in O(f)$. The ReLU activation function is given for every $x\in\mathbb{R}$ by $\text{ReLU}(x)=\max\{x,0\}$. For each $n\in\mathbb{N}_{+}$ and $C\subseteq\mathbb{R}^{n}$, the indicator function $I_{C}$ of $C$ is defined by: for each $x\in\mathbb{R}^{n}$ set $I_{C}(x)=1$ if $x\in C$ and is 00 otherwise.
Multi-Layer Perceptrons. In this paper, we will consider MLPs with trainable PReLU activation functions.
Definition 3.1 (Trainable PReLU).
We define the trainable PReLU activation function $\sigma:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ for each input $x\in\mathbb{R}$ and each parameter $\alpha\in\mathbb{R}$ as follows:
$\sigma_{\alpha}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sigma(x,\alpha)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\begin{cases}x~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{ if }x\geq 0,\\ \alpha x~{}~{}~{}~{}~{}~{}\text{ otherwise.}\end{cases}$
PReLU generalizes ReLU since $\text{ReLU}(x)=\sigma_{0}(x)$, and it makes the hyperparameter $\alpha$ of a Leaky ReLU learnable. We will often be applying our trainable activation functions component-wise. For positive integers $n,m$, we denote the set of $n\times m$ matrices by $\mathbb{R}^{n\times m}$. More precisely, we mean the following operation defined for any $N\in\mathbb{N}$, $\bar{\alpha}\in\mathbb{R}^{N}$ with $i^{th}$ entry denoted as $\bar{\alpha}_{i}$, and $x\in\mathbb{R}^{N}$, by
$\sigma_{\bar{\alpha}}\bullet x\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\left(\sigma_{\bar{\alpha}_{i}}(x_{i})\right)_{i=1}^{N}.$
We now define the class of multilayer perceptions (MLPs), with trainable activation functions. Fix $J\in\mathbb{N}$ and a multi-index $[d]\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(d_{0},\dots,d_{J+1})$, and let $P([d])=\sum_{j=0}^{J}d_{j}(d_{j+1}+2)$. We identify any vector $\theta\in\mathbb{R}^{P([d])}$ with
$\theta$ $\leftrightarrow\big{(}A^{(j)},b^{(j)},\bar{\alpha}^{(j)}\big{)}_{j=0}^{J}$ (3)
$(A^{(j)},b^{(j)},\bar{\alpha}^{(j)})$ $\in\mathbb{R}^{d_{j+1}\times d_{j}}\times\mathbb{R}^{d_{j}}\times\mathbb{R}^{d_{j}}$
With the above identification, we recursively define the representation function of a $[d]$-dimensional network by
$\mathbb{R}^{P([d])}\times\mathbb{R}^{d_{0}}\ni(\theta,x)$ $\mapsto\hat{f}_{\theta}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}A^{(J)}\,x^{(J)}+b^{(J)},$ (4)
$x^{(j+1)}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sigma_{\bar{\alpha}^{(j)}}\bullet(A^{(j)}\,x^{(j)}+b^{(j)})$
$x^{(0)}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}x.$
where $j$ runs from 00 to $J-1$.
We denote by $\mathcal{NN}_{[d]}^{\sigma}$ the family of $[d]$-dimensional multilayer perceptrons (MLPs), $\{\hat{f}_{\theta}\}_{\theta\in\mathbb{R}^{P([d])}}$ described by (4). The subset of $\mathcal{NN}_{[d]}^{\sigma}$ consisting of networks $\hat{f}_{\theta}$ with each $\bar{\alpha}^{(j)}_{i}=(1,0)$ in Eq. (4) is denoted by $\mathcal{NN}_{[d]}^{\operatorname{ReLU}}$ and consists of the familiar deep ReLU MLPs. The set of ReLU MLPs with depth $J$ and width $W$ is denoted by $\mathcal{NN}_{J,W:n,m}^{\sigma}=\cup_{[d]}\,\mathcal{NN}_{[d]}^{\sigma}$, where the union is taken over all multi-indices $[d]=[d_{0},\dots,d_{\tilde{J}}]$ with $n=d_{0}$, $m=d_{J+1}$, $d_{0},\dots,d_{J+1}\leq W$, and $\tilde{J}\leq J$.
We now formally define the classes of neural pathways.
Definition 3.2 (Neural Pathways).
Let $J,W,L,n\in\mathbb{N}$ and fix an activation function $\sigma\in C(\mathbb{R})$. The set of neural pathways with at-most $L$ leaves, depth $J$, and width $W$, denoted by $\mathcal{NP}_{J,W,L:n,m}^{\sigma}$, consists of all functions $\hat{f}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ satisfying $\hat{f}=\sum_{i=1}^{l}\,f_{i}\,I_{C_{i}}$ where for $f_{1},\dots,f_{l}\in\mathcal{NN}^{\operatorname{PReLU}}_{J,W}$, and distinct prototypes $p_{1},\dots,p_{L}\in\mathbb{R}^{n}$; whose induced Voronoi cells $\{C_{i}\}_{i=1}^{L}$ are defined recursively by
$C_{i}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{C}_{i}\setminus\bigcup_{j<i}\,\tilde{C_{j}}$ (5)
where for $i=1,\dots,l$ the (non-disjoint) cells are
$\tilde{C}_{i}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{\{}x\in[0,1]^{n}:\,\|x-p_{i}\|=\min_{j\in\{1,\dots,l\}}\,\|x-p_{j}\|\big{\}}$
Remark 3.3 (Trees).
The structure in any neural pathway is any tree with root node $\mathbb{R}^{n}$ and leaves given by the pairs $\{(p_{i},f_{i})\}_{i=1}^{l}$ or equivalently $\{(C_{i},f_{i})\}_{i=1}^{l}$. The purpose of any such tree is simply to efficiently route an input $x\in\mathbb{R}^{n}$ to one of the $l$ “leaf networks” $f_{1},\dots,f_{l}$ using $\mathcal{O}(l\log(l))$ queries; to identify which Voronoi cells $\{C_{i}\}_{i=1}^{l}$ the point $x$ is contained in.
The partitioning technique used to define Eq. (5) is standard, see e.g. Krauthgamer et al. (2005, Proof of Lemma 1.7; page 846). It is employed to ensure disjointness of the Voronoi cells; this guarantees that no input is assigned to multiple prototypes. To keep notation tidy, we use $\mathcal{NP}_{J,W,L}^{\sigma}$ (resp. $\mathcal{N}^{\sigma}_{W,L}$), to abbreviate $\mathcal{NP}_{J,W,L:n,m}^{\sigma}$ (resp. $\mathcal{N}^{\sigma}_{W,L:n,m}$) whenever $n$ and $m$ are clear from the context.
VC dimension
Let $\mathcal{F}$ be a set of functions from a subset $\mathcal{X}\subseteq\mathbb{R}^{n}$ to $\{0,1\}$; i.e. binary classifiers on $\mathcal{X}$. The set $\mathcal{F}$ shatters (in the classical sense) a $k$-point subset $\{x_{i}\}_{i=1}^{k}\subseteq\mathcal{X}$ if $\mathcal{F}$ can represent every possible set of labels on those $k$-points; i.e. if #$\#\,\{(\hat{f}(x_{i}))_{i=1}^{k}\in\{0,1\}^{k}:\,\hat{f}\in\mathcal{F}\}=2^{k}$.
As in Shen et al. (2022b), we extend the definition of shattering from binary classifiers to real-valued functions as follows. Let $\mathcal{F}$ be a set of functions from $[0,1]^{n}$ to $\mathbb{R}$. The set $\mathcal{F}$ is said to shatter a $k$-point set $\{x_{i}\}_{i=1}^{k}\subseteq\mathcal{X}$ if
$\{I_{(0,\infty)}\circ f:\,f\in\mathcal{F}\}$ (6)
shatters it in the classical sense; here $I_{(0,\infty)}(t)=1$ if $t\geq 0$ and equals to 00 otherwise. Denoted by $\operatorname{VC}(\mathcal{F})$, the VC dimension of $\mathcal{F}$ is the cardinality of the largest $k$-point subset shattered by $\mathcal{F}$. If $k$ is unbounded, then we say that $\mathcal{F}$ has an infinite VC dimension (over $\mathcal{X}$). One can show, see Bartlett et al. (2019), that the VC-dimension of any such $\mathcal{F}$ is roughly the same as the pseudo-dimension of Pollard (1990) for a small modification of $\mathcal{F}$.
VC dimension measures the richness of a class of functions. For example, in Harvey et al. (2017, Theorem 1), the authors showed that the set of MLPs with $\operatorname{ReLU}$ activation function with $L\in\mathbb{N}_{+}$ layers, width and $W\in\mathbb{N}_{+}$ satisfying $W>O(L)>C^{2}$, where $C\geq 640$, satisfies
$\operatorname{VC}(\mathcal{NN}_{W,L}^{\sigma})\in\Omega\Big{(}WL\,\log_{2}(W/L)\Big{)}.$ (7)
Nearly matching upper bounds are in Bartlett et al. (2019).
3.2 Main Theoretical Results
To reduce notation, our main result is formulated for locally-Hölder function on the closed unit ball of $\mathbb{R}^{n}$, defined by $\overline{B}_{n}(0,1)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{x\in\mathbb{R}^{n}:\,\|x\|\leq 1\}$.
Theorem 3.4 (Neural Pathways Avoid The Curse of Dimensionality).
Fix $0<\alpha\leq 1$, dimensions $n,m\in\mathbb{N}$, parameters $0<\alpha\leq 1$ and an $\alpha$-Hölder map $f:\overline{\operatorname{B}}_{n}(0,1)\to\mathbb{R}^{m}$ with optimal Hölder coefficient $L\geq 0$, and $\sigma\in C(\mathbb{R})$ satisfying Definition 3.1.
Fix an “approximation error” $\varepsilon>0$. There is a binary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq\overline{\operatorname{B}}_{n}(0,1)\times\mathbb{R}^{P([d])}$ satisfying
$\max_{x\in\overline{\operatorname{B}}_{n}(0,1)}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|\in\mathcal{O}\big{(}\varepsilon^{1/\alpha-1/n}\big{)}.$ (8)
Furthermore, for each $x\in\overline{\operatorname{B}}_{n}(0,1)$, if $v_{i}$ is such that $\|x-v_{i}\|\in\mathcal{O}(\varepsilon^{1/\alpha-1/n})$ then
$\|f(x)-\hat{f}_{\theta_{i}}(x)\|<\varepsilon.$ (9)
The depth and width of each $\hat{f}_{\theta_{i}}$ are recorded in Table 1 and the binary tree $\mathcal{T}$ has $2^{\mathcal{O}(n-\frac{n}{\alpha}\log(\varepsilon))}$  leaves.
Remark 3.5.
A technical version of Theorem 3.4 with $\nu$-ary trees and general uniformly continuous target functions, namely Theorem A.5, can be found in Appendix A.2.
Next, we demonstrate that the neural pathways model can generalize and generate functions that are PAC-learnable, thanks to its finite VC dimension. This property, however, breaks down in MLP models with super-expressive activation functions.
Proposition 3.6 (Neural Pathways Can Generalize).
Let $J,W,L,n\in\mathbb{N}_{+}$. Then $\operatorname{VC}\big{(}\mathcal{NN}_{J,W,L:n,1}^{\operatorname{PReLU}}\big{)}$ is of
$\mathcal{O}\big{(}L\log(L)^{2}\,\max\{nL\log(L),JW^{2}\log(JW)\}\big{)}$
In particular, $\mathrm{VC}(\mathcal{NP}_{J,W,L:n,1}^{ReLU})<\infty$.
We complement the main result of Bartlett et al. (2019) by demonstrating that the class of unstable MLPs (Shen et al., 2022a) possesses infinite VC dimension, even when they have finite depth and width. Thus, while they may serve as a gold standard from the perspective of approximation theory, they should not be considered a benchmark gold standard from the viewpoint of learning theory.
Proposition 3.7 (MLPs with Super-Expressive Activation Do Not Generalize).
Let $\mathcal{F}$ denote the set of MLPs with activation function in Definition A.1, depth at-most $15$, and width at-most $36n(2n+2)$. Then $\operatorname{VC}(\mathcal{F})=\infty$.
The VC dimension bounds for the standard MLP model, MLP with a super-expressive activation function as proposed by Shen et al. (2022a), and the neural pathways model are summarized in Table 2.
Table 2: VC Dimension of the Neural Pathways, ReLU MLP, and MLP model with Super-Expressive Activation function of Shen et al. (2022a). All models have depth $J$, width $W$, and (when applicable) $L$ leaves; where $J,W,L,n\in\mathbb{N}_{+}$.
Model VC Dim Ref. Neural Pathways $\mathcal{O}\big{(}L\log(L)^{2}\,\max\{nL\log(L),JW^{2}\log(JW)\}\big{)}$ Prop 3.6 ReLU MLP $\mathcal{O}\big{(}JW^{2}\log(JW)\big{)}$ Bartlett et al. (2019) Super-Expressive $\infty$ Prop 3.7
4 Related Work and Discussion
With the advent of large foundation models, the use of distributed computing to scale deep learning models beyond what is feasible to process on a single machine has become exceedingly relevant. However, most of the literature focuses on distributed or parallelized GPU training of standard neural network models (Dean et al., 2012; Hegde & Usmani, 2016; Sergeev & Del Balso, 2018; Ben-Nun & Hoefler, 2019). Next, we review relevant literature.
Prototypes. Each region in our learned partition of the input space is associated with a representative point therein called a prototype. Prototypes (also called landmarks) are routinely used in image classification (Mensink et al., 2012), few-shot learning (Snell et al., 2017; Cao et al., 2020), dimensionality reduction (Law et al., 2019), in complex networks (Keller-Ressel & Nargang, 2023), and geometric deep learning (Ghadimi Atigh et al., 2021) to tractably encode massive structures. They are also standard in classical clustering algorithms such as $K$-medoids or $K$-means, wherein the part associated with each medoid (resp. centroid) defines a Voronoi cell or Voronoi region (Voronoi, 1908). In our proposed approach, one key difference is that we optimize prototypes to maximize performance on the loss function defining the learning problem, rather than to divide up the input space based on a clustering objective, which may not optimize the model’s performance. Nevertheless, in principle, many of these prototype-based clustering algorithms can be substituted into our pipeline.
Partitioning. While partitioning is commonly employed in deep learning for various purposes, such as proving universal approximation theorems (Yarotsky, 2017; Lu et al., 2021b; Gühring & Raslan, 2021) or facilitating clustering-based learning (Kratsios & Zamanlooy, 2022; Trask et al., 2022; Ali & Nouy, 2023; Srivastava et al., 2022), existing approaches typically involve loading the entire model into VRAM. Our approach, however, differs by relying on a learned partition of the input space, where each part is associated with a distinct small neural network. Importantly, the complete set of networks forming the neural pathways does not need to be simultaneously loaded into VRAM during training or inference.
Cluster-then-Predict. Several deep learning models employ a cluster-then-predict approach to upscaling, e.g., (Wang et al., 2021). In these approaches, one typically maps an input to a specific prototype and passes that prototype through a relatively small network, effectively reducing dimension by discretizing the input space to consist only of the set of prototypes. This is in stark contrast to the neural pathways framework, which performs no discretization and only hierarchically partitions the input space. Every input is passed through a neural network at the leaf node, corresponding to the nearest prototypes, but the input itself is never replaced by one of the prototypes. Srivastava et al. (2022) also use multiple local networks, but they employ soft clustering for the final prediction, which means they have to load all the local models for a single sample.
Comparison to Standard Distributed Training. One can distribute the complexity of feedforward models by storing each of their layers in offline memory and then loading them sequentially into VRAM during the forward pass. This does avoid loading more than $\mathcal{O}(\operatorname{Width})$ active parameters into VRAM at any given time, where $\operatorname{Width}$ denotes the width of the feedforward model. However, doing so implies that all the model parameters are ultimately loaded during the forward pass. This contrasts with the neural pathways model, which requires $\mathcal{O}\big{(}\log_{2}(K)\,\operatorname{Width}^{2}\,\operatorname{Depth}\big{)}$ to be loaded into memory during a forward pass; where $\operatorname{Width}$ and $\operatorname{Depth}$ are respectively the largest width and depth of the MLP at any leaf of the tree defining a given neural pathway, and $K$ denotes the number of prototypes. However, in the forward pass, one loads $\mathcal{O}(\varepsilon^{-n/2})$ parameters for the best worst-case MLP while only $\mathcal{O}(n\log(1/\varepsilon)/\varepsilon)$ are needed in the case of the neural pathways. The number of parameters here represents the optimal worst-case rates for both models (see Table 1 and Theorem A.5).
Pruning. Additionally, one might consider the option of pruning a sizable model, conceivably trained on a GPU with a larger VRAM, for utilization on a smaller GPU during inference as an alternative to our method. Nevertheless, in frameworks like PyTorch, pruning does not result in a reduction of operations, acceleration, or diminished VRAM memory usage. Instead, pruning only masks the original model weights with zeros. The reduction in model size occurs only when saved in offline memory in sparse mode, which, in any case, is not a significant concern.
Logarithmic number of queries via trees. For many prototypes, as in our main guarantee, the neural pathways only need to evaluate the distance between the given input and a logarithmic number of prototypes—specifically, one for each level in the tree—when using deep binary trees to hierarchically refine the Voronoi cells. Thus, a given machine never processes the exponential number of prototypes, and only $\nu\lceil\log_{\nu}(K)\rceil$ prototypes are ever queried for any given input; when trees are $\nu$-ary (as in Theorem A.5), and where $K$ denotes the number of prototypes. Since we consider that prototypes are queried separately and before loading neural pathways, we do not take them into account when counting the number of learnable parameters. Moreover, the size of our prototypes is negligible in our experiments.
5 Experiments
We evaluate our approach in two standard machine learning tasks: regression and classification. We experimentally show that neural pathways, which distribute predictions over multiple neural networks, are competitive with a single large neural network containing as many model parameters as all the neural pathways combined. This is desirable in cases where the large neural network does not fit into the memory of a single machine. On the contrary, the neural pathways framework can be trained by distributing each neural pathway on separate machines (or equivalently serially on a single machine). Inference can then be performed by loading only a single neural pathway at a time into the GPU. Our source code to reproduce our results is available in the supplementary material.
5.1 Regression
We first consider regression, where the goal is to approximate non-convex synthetic functions often used for performance test problems. In particular, we choose 1-dimensional Hölder functions, as well as Ackley (Ackley, 1987) and Rastrigin (Rastrigin, 1974) functions, whose formulations are detailed in Appendix D.
1D Hölder Functions. We illustrate our primary finding, as encapsulated in Theorem 3.4, by leveraging $1$-dimensional functions characterized by very low regularity. This choice is motivated by the jagged structure inherent in such functions, necessitating an exponentially higher sampling frequency compared to smooth functions for achieving an accurate reconstruction. Indeed, this crucial sampling step forms the foundation of many quantitative universal approximation theorems (Yarotsky, 2017; Shen et al., 2021; Kratsios & Papon, 2022). As elaborated in Appendix C, approximating a well-behaved (Lipschitz) function in $d$ dimensions poses a challenge equivalent to approximating a highly irregular function ($1/d$-Hölder) in a single dimension. A visual representation of a $1/d$-Hölder function is presented in Figure 2, exemplified by the trajectory of a fractional Brownian motion with a Hurst parameter of $\alpha=1/d$. A formal definition is available in Appendix C.
Figure 2: Visual Comparison of Functions with High ($\alpha\approx 1$) vs. Low ($\alpha\approx 0$) Hölder regularity. If $\alpha\approx 1$, the function (green) is approximately differentiable almost everywhere, meaning it does not osculate much locally and thus is simple to approximate. If $\alpha\approx 0$, the function may be nowhere differentiable and jagged; its extreme details make it difficult to approximate.
Table 3: Test mean squared error (average and standard deviation) for the different functions of the regression task.
1D Hölder 2D Ackley 3D Ackley 2D Rastrigin 3D Rastrigin
Neural pathways (ours) 0.057 $\pm$ 0.085 0.00015 $\pm$ 0.00006 0.00068 $\pm$ 0.00010 0.0480 $\pm$ 0.0073 1.0062 $\pm$ 0.0446
Baseline 0.128 $\pm$ 0.012 0.08723 $\pm$ 0.01059 0.09303 $\pm$ 0.03156 3.0511 $\pm$ 0.3581 8.0376 $\pm$ 4.0499
2D and 3D Functions. We select the Ackley and Rastrigin functions, with their respective 2D representations showcased in Figure 3, as widely recognized benchmarks in the field of optimization.
(a) Ground Truth Ackley
(b) Ground Truth Rastrigin
(c) Predicted Ackley
(d) Predicted Rastrigin
(e) log MSE
(f) log MSE
Figure 3: Comparison of ground truth and predicted results for 2D Ackley and Rastrigin functions over the domain $[-1,1]^{2}$.
Evaluation protocol. We consider the setup where the domain of a function that we try to approximate is the $n$-dimensional closed set $[a,b]^{n}$. For instance, we arbitrarily choose the domain $[0,1]$ when $n=1$, and $[-1,1]^{n}$ when $n\geq 2$. Our training and test samples are the $s^{n}$ vertices of the regular grid defined on $[a,b]^{n}$. At each run, $80\%$ of the samples are randomly selected for training and validation, and the remaining $20\%$ for testing.
During training, for a given and fixed set of $K$ prototypes $p\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(p_{1},\dots,p_{K})$, we assign each training sample $x$ to its nearest prototype $p_{k}$ and associated neural network $\hat{f}_{k}$. We learn the prototypes as explained in Section 2.2. For simplicity, we set the number of prototypes to $K=4$; we also set $s=10,000$ if $n=1$, $s=150$ if $n=2$ (i.e., $500150^{2}=22,$ 500 samples), and $s=30$ if $n=3$ (i.e., $27,000$ samples). More details can be found in Appendix D.
Test performance. In Table 3, we present the mean squared error obtained on the test set across 10 random initializations and various splits of the training/test sets. The baseline consists of a single neural network with the same overall architecture as each neural pathway but possesses as many parameters as all the neural pathways combined. In all instances, the neural pathways model demonstrates a significant performance advantage compared to the baseline. Figure 3 illustrates the predictions generated by our neural pathways, showcasing the capability of our approach to achieve a good approximation of the ground truth functions.
5.2 Classification
Datasets. We evaluate classification on standard image datasets such as CIFAR-10 (Krizhevsky & Hinton, 2010), CIFAR-100, and Food-101 (Bossard et al., 2014), which consist of 10, 100, and 101 different classes, respectively. We use the standard splits of training/test sets: the datasets include (per category) 5,000 training and 1,000 test images for CIFAR-10, 500 training and 100 test images for CIFAR-100, and 750 and 250 for Food-101.
Training. Our neural pathways model takes as input latent DINOv2 encodings (Oquab et al., 2023) of images from the aforementioned datasets. Each sample $x\in\mathbb{R}^{768}$ corresponds to a DINOv2 embedding (i.e., $n=768$). Additionally, we set the prototypes as centroids obtained through the standard $K$-means clustering on the DINOv2 embedding space. The replacement of our original prototype learning algorithm is sensible in this context, as we operate within a structured latent space optimized through self-supervised learning using large-scale compute and datasets.
Table 4: Test classification accuracy using DINOv2 features as input (average and standard deviation).
Dataset CIFAR-10 CIFAR-100 Food-101
Ours (weighted) 98.40 $\pm$ 0.05 90.01 $\pm$ 0.11 91.86 $\pm$ 0.10
Ours (unweighted) 98.42 $\pm$ 0.04 89.62 $\pm$ 0.25 91.79 $\pm$ 0.16
Baseline 98.45 $\pm$ 0.06 89.85 $\pm$ 0.17 91.45 $\pm$ 1.09
Due to the potential class imbalance in the various Voronoi cells formed by the prototypes, we utilize two variations of the cross-entropy loss for each neural pathway $\hat{f}_{k}$. The first variation, termed unweighted, assigns equal weight to all categories. The weighted variation assigns a weight that is inversely proportional to the distribution of each category in the Voronoi cell defined by the prototype $p_{k}$.
Test performance. We present the test classification accuracy over 10 random initializations of both the baseline and our neural pathways in Table 4. The weighted version performs slightly better on datasets with a large number of categories. Our approach achieves comparable results with the baseline, effectively decomposing the prediction across multiple smaller models.
6 Conclusion
We introduced a novel approach to overcome the curse of dimensionality based on distributed neural computation. Instead of relying on exceptionally regular target functions or super-expressive MLPs, as seen in other related works in the literature, we propose distributing computation across several small specialized networks. These networks can learn to model the target function within Voronoi cells of the input space. Our approach is particularly relevant, given the computational limitations of current AI hardware, handicapped by finite VRAM. This problem has recently been exacerbated by the growing trend to increase model size with large foundation models. In particular, our neural pathways framework can scale up simply by adding more neural networks (or equivalently by refining the Voronoi cells which subdivide the input space), since each independent network can always be loaded into VRAM while the overall model keeps increasing in capacity.
We prove, via Theorem 3.4 and A.5, that the number of parameters of any single neural pathway loaded in active GPU memory only requires $\mathcal{O}(n/\varepsilon)$ depth and $\mathcal{O}(n)$ width when approximating a Lipschitz map over the cube $[0,1]^{n}$ uniformly to an accuracy of $\varepsilon>0$. This contrasts with the optimal approximation rates for standard (non-distributed) MLP architectures, which require $\mathcal{O}(\varepsilon^{-n})$ depth and $\mathcal{O}(n)$ width to achieve the same $\varepsilon$ accuracy.
Importantly, unlike results that break the curse of dimensionality using standard MLPs, our results always hold in practice since we do not require the target function to have high regularity. Furthermore, in contrast to approximation results leveraging super-expressive activation functions with infinite VC-dimension (Proposition 3.7), any neural pathway of finite depth and width, defined by a finite tree with weight and bias constraints, has finite VC-dimension (Proposition 3.6). Thus, our models can learn and generalize effectively when trained on finite data.
We validate our algorithmic framework and mathematical analysis by testing neural pathways on regression tasks. Specifically, we focus on learning 1D Hölder functions of low regularity, as well as on 2D and 3D Ackley and Rastrigin functions. Additionally, we assess its performance in classification, particularly in higher-dimensional spaces such as the latent space of pretrained self-supervised large computer vision models like DINOv2. This not only helps verify the applicability of our model beyond synthetic experiments but also demonstrates its effectiveness in setups that are of closer interest to the applied machine learning community. In all cases, our model either matches or outperforms the centralized baseline while loading substantially lighter networks into GPU memory one at a time.
Future Work. In this paper, we present the initial implementation of our framework and we anticipate that future efforts will be necessary to scale it to larger networks. In such cases, the inclusion of deeper tree structures, akin to those discussed in our proofs, will be pertinent in addressing computational complexity. Similarly, as discussed in Section 2, our framework is modular, and we foresee that the methodology could be potentially modified depending on the downstream task and specific network architectures, as we primarily focused on MLPs only.
7 Acknowledgment and Funding
The authors would like to thank James Lucas and Rafid Mahmood for their helpful feedback while preparing the manuscript.
AK acknowledges financial support from the NSERC Discovery Grant No. RGPIN-2023-04482. HSOB acknowledges financial support from an EV travel grant.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Overview of Technical Derivation
We now overview the proof of our main result and its full technical formulation. These objectives require us to recall definitions from the analysis of metric spaces, which we do now.
A.1 Preliminary Definitions
The metric ball in $(\mathcal{X},d)$ of radius $r>0$ at $x\in\mathcal{X}$ is denoted by $\operatorname{Ball}_{(\mathcal{X},d)}(x,r)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{z\in\mathcal{X}\colon d(x,z)<r\}$.
The metric space $(\mathcal{X},d)$ is called doubling, if there is $C\in\mathbb{N}_{+}$ for which every metric ball in $(\mathcal{X},d)$ can be covered by at most $C$ metric balls of half its radius. The smallest such constant is called $(\mathcal{X},d)$’s doubling number, and is here denoted by $C_{(\mathcal{X},d)}$. Though this definition may seem abstract at first, (Heinonen, 2001, Theorem 12.1), provides an almost familiar characterization of all doubling metric spaces; indeed, $\mathcal{K}$ is a doubling metric space if and only if it can be identified via a suitable invertible map [^1] So called quasi-symmetric maps, see (Heinonen, 2001, page 78). with a subset of some Euclidean space. Every subset of $\mathbb{R}^{n}$, for any $n\in\mathbb{N}_{+}$, is a doubling metric space; see (Robinson, 2011, Chapter 9) for details.
Example 1 (Subsets of Euclidean Spaces).
Fix a dimension $n\in\mathbb{N}_{+}$. The doubling number of any subset of Euclidean space is [^2] See (Robinson, 2011, Lemma 9.2) and the brief computations made in the proof of (Robinson, 2011, Lemma 9.4). at most $2^{n+1}$.
In what follows, all logarithms will be taken base $2$, unless explicitly stated otherwise, i.e. $\log_{v}$ is base $v$ for a given $v\in\mathbb{N}_{+}$ and $\log\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log_{2}$. As on (Petrova & Wojtaszczyk, 2023, page 762), the radius of a subset $A\subseteq\mathbb{R}^{n}$, denoted by $\operatorname{rad}(A)$, is defined by
$\operatorname{rad}(A)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\inf_{x\in\mathbb{R}^{n}}\,\sup_{a\in A}\,\|x-a\|.$
The diameter of any such set $A$, denoted by $\operatorname{diam}(A)$, satisfies the sharp inequality $\operatorname{diam}(A)\leq 2\operatorname{rad}(A)$.
Finally, let us recall the notion of a uniformly continuous function. Fix $n,m\in\mathbb{N}_{+}$ and let $X\subset\mathbb{R}^{n}$. Let $\omega:[0,\infty)\to[0,\infty)$ be a monotonically increasing function which is continuous at 00 and satisfies $\omega(0)=0$. Such an $\omega$ is called a modulus of continuity. A function $f:X\rightarrow\mathbb{R}^{m}$ is said to be $\omega$-uniformly continuous if
$\|f(x)-f(y)\|\leq\omega\big{(}\|x-y\|\big{)}$
holds for all $x,y\in\mathcal{X}$. We note that every continuous function is uniformly continuous if $\mathcal{X}$ is compact and that its modulus of continuity may depend on $\mathcal{X}$. Furthermore, we note that every $(\alpha,L)$-Hölder function is uniformly continuous with modulus of continuity $\omega(t)=L\,t^{\alpha}$.
We consider a mild extension of the super-expressive activation function of Shen et al. (2022a). This parametric extension allows it to implement the identity map on the real line as well as the original super-expressive activation function thereof.
Definition A.1 (Trainable Super-Expressive Activation Function).
A trainable action function $\sigma:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is of super-expressive type if for all $\alpha\in\mathbb{R}$
$\sigma_{\alpha}:\mathbb{R}\ni x\mapsto\alpha x+(1-\alpha)\sigma^{\star}(x)$
where $\sigma^{\star}:\mathbb{R}\to\mathbb{R}$ is given by
$\sigma^{\star}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\begin{cases}|x\mod{2}|&\mbox{ for }0\leq x\\ \frac{x}{|x|+1}&\mbox{ for }x<0.\end{cases}$ (10)
A.2 Proof Sketch
Lemma A.2.
Let $\mathcal{K}$ be a compact subset of $\mathbb{R}^{n}$ whose doubling number is $C$. Fix $v\in\mathbb{N}$ with $v\geq 2$, and $0<\delta\leq\operatorname{rad}(\mathcal{K})$. There exists an $v$-ary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves $\mathcal{L}\subseteq\mathcal{K}$ satisfying
$\max_{x\in\mathcal{K}}\,\min_{v\in\mathcal{L}}\,\|x-v\|<\delta.$ (11)
Furthermore, the number of leaves $L\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\#\mathcal{L}$, height, and total number of nodes #$\#V$ of the tree $\mathcal{T}$ are
  • (i)
    Leaves: at most $L=v^{\big{\lceil}c\,\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil},}$
  • (ii)
    Height: $\big{\lceil}c\,\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil}$,
  • (iii)
    Nodes: At most $\frac{v^{\lceil c\,\log(C)(1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}})))\rceil+1}-1}{\big{\lceil}c\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil}-1}$
where $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}1/\log(v)$.
At each node of the tree, we will place an MLP which only locally approximates the target function on a little ball of suitable radius (implied by the tree valency $v$ and height $h$) of lemma A.2. I.e. by the storage space we would like to allocate to our neural pathway model. The next step of the proof relies on a mild extension of the quantitative universal approximation theorem in Shen et al. (2022a); Lu et al. (2021a) to the multivariate case, as well as an extension of the multivariate approximation result of Acciaio et al. (2023, Proposition 3.10) beyond the Hölder case.
Lemma A.3 (Universal Approximation Theorem).
Let $n,m\in\mathbb{N}_{+}$ with $n\geq 3$, ${\mathcal{K}}\subseteq\mathbb{R}^{n}$ be compact set $f:{\mathcal{K}}\to\mathbb{R}^{m}$ be uniformly continuous with strictly monotone continuous modulus of continuity $\omega$. Let $\sigma$ be an activation function as in Definitions A.1 or 3.1. For each $\varepsilon>0$, there exists an MLP $\hat{f}_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with trainable activation function $\sigma$ satisfying the uniform estimate
$\sup_{x\in{\mathcal{K}}}\,\|f(x)-\hat{f}_{\theta}(x)\|\leq\epsilon.$
The depth and width of $\hat{f}$ are recorded in Table 5.
Table 5: Complexity of the MLP $\hat{f}_{\theta}$ in Lemma A.3.
Activation $\sigma$ Super Expressive A.1 PReLU 3.1
Depth $(J)$ $15m$ $m\left(19+2n+11\Big{\lceil}\Big{(}\frac{\operatorname{rad}(K)2^{3/2}n^{1/2}}{(n+1)^{1/2}\omega^{-1}\big{(}\varepsilon/(131\sqrt{n\,m})\big{)}}\Big{)}^{n/2}\Big{\rceil}\right)$
Width ($\max_{j}\,d_{j}$) $36n(2n+1)+m$ $16\max\{n,3\}+m$
Combining Lemmata A.2 and A.3 we obtain Theorem 3.4.
Lemma A.4.
Let ${\mathcal{K}}$ be a compact subset of $\mathbb{R}^{n}$ whose doubling number is $C$ and a uniformly continuous map $f:{\mathcal{K}}\to\mathbb{R}^{m}$ with modulus of continuity $\omega$.
Fix $v\in\mathbb{N}$ with $v\geq 2$, $0<\delta\leq\operatorname{rad}({\mathcal{K}})$, and $\varepsilon>0$. Suppose that $\sigma$ satisfies Definition 3.1. There exists a $p\in\mathbb{N}_{+}$ and a $v$-ary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq{\mathcal{K}}\times\mathbb{R}^{p}$ satisfying
$\max_{x\in{\mathcal{K}}}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|<\delta.$ (12)
Furthermore, for each $x\in{\mathcal{K}}$ and each $i=1,\dots,L$, if $\|x-v_{i}\|<\delta$ then
$\|f(x)-f_{\theta_{i}}(x)\|<\varepsilon.$
We have the following estimates:
  1. (i)
    Depth: Depth of each $\hat{f}_{\theta_{i}}$ is $m\left(19+2n+11\Big{\lceil}\Big{(}\frac{\delta 2^{3/2}n^{1/2}}{(n+1)^{1/2}\omega^{-1}\big{(}\varepsilon/(131\sqrt{n\,m})\big{)}}\Big{)}^{n/2}\Big{\rceil}\right)$
  2. (ii)
    Width: Width of each $\hat{f}_{\theta_{i}}$ is $16\,\max\{n,3\}+m$
  3. (iii)
    Leaves: at most $L=v^{\big{\lceil}c\,\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil},}$
  4. (iv)
    Height: $\big{\lceil}c\,\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil}$,
  5. (v)
    Nodes: At most $\frac{v^{\lceil c\,\log(C)(1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}})))\rceil+1}-1}{\big{\lceil}c\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil}-1}$
where $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
We now present the technical version of Theorem 3.4. This result allows distributed neural computing using $\nu$-ary trees and allows for the approximation general uniformly continuous target functions.
Theorem A.5 (Technical Version of Theorem 3.4).
Let ${\mathcal{K}}$ be a compact subset of $\mathbb{R}^{n}$ whose doubling number is $C$ and a uniformly continuous map $f:{\mathcal{K}}\to\mathbb{R}^{m}$ with modulus of continuity $\omega$. Fix an “approximation rate” $r\geq 0$, a “valency parameter” $v\in\mathbb{N}$ with $v\geq 2$, and an “approximation error” $\varepsilon>0$. Suppose that $\sigma$ satisfies Definition 3.1. Then, there is an $v$-ary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq{\mathcal{K}}\times\mathbb{R}^{p}$ satisfying
$\max_{x\in{\mathcal{K}}}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|<\frac{\varepsilon^{-2r/n}}{2}\,\omega^{-1}\biggl{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\biggr{)}.$
Furthermore, for each $x\in{\mathcal{K}}$ and $i=1,\dots,L$, if $\|x-v_{i}\|<\delta$ then
$\|f(x)-f_{\theta_{i}}(x)\|<\varepsilon.$
The depth and width of each $\hat{f}_{\theta_{i}}$ and the number of leaves, height, and number of nodes required to build the $\nu$-ary tree are all recorded in Table 6.
Table 6: Complexity of Feedforward Neural Network $\hat{f}_{\theta}$ and the $\nu$-ary tree in Theorem A.5. Here $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
Parameter Estimate
Depth $(J)$ $m\left(19+2n+11\lceil\varepsilon^{-r}\rceil\right)$
Width ($\max_{j}\,d_{j}$) $16\max\{n,3\}+m$
Leaves $\mathcal{O}\Big{(}v^{\big{\lceil}c\,\log(C)\big{(}1+\log(\varepsilon^{2r/n}\,\operatorname{diam}({\mathcal{K}}))/\big{(}2\,\omega^{-1}\big{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\big{)}\big{)}\big{)}\big{\rceil},}\Big{)}$
Height $\big{\lceil}c\,\log(C)\big{(}1+\log(\epsilon^{2r/n}\operatorname{diam}({\mathcal{K}})/\big{(}2\omega^{-1}\big{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\big{)}\big{)})\big{)}\big{\rceil}$
Nodes $\mathcal{O}\Biggl{(}\frac{v^{\lceil c\,\log(C)(1+\log(\varepsilon^{2r/n}\operatorname{diam}({\mathcal{K}})/\big{(}2\omega^{-1}(\frac{\varepsilon}{131\,(nm)^{1/2}})\big{)}))\rceil+1}-1}{\big{\lceil}c\log(C)\big{(}1+\log(\frac{\varepsilon^{2r/n}}{2}\operatorname{diam}({\mathcal{K}}))/\omega^{-1}\biggl{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\biggr{)}\big{)}\big{\rceil}-1}\Biggr{)}$
Appendix B Appendix: Proof of Main Theorem
Proof of Lemma A.2.
Since ${\mathcal{K}}$ is a doubling metric space then, (Acciaio et al., 2023, Lemma 7.1), for each $\delta>0$, there exist $x_{1},\dots,x_{N}\in\mathcal{K}$ satisfying
$\max_{x\in\mathcal{K}}\,\min_{i=1,\dots,N_{\delta}}\,\|x-x_{i}\|<\delta\mbox{ and }N_{\delta}\leq C^{\lceil\log(\operatorname{diam}(\mathcal{K})/\delta)\rceil}.$
In particular, since the doubling number of $\mathcal{K}$ is $C$, we have the upper-bound
$N_{\delta}\leq C\,C^{\log(\delta^{-1}\operatorname{diam}(\mathcal{K}))}.$ (13)
An elementary computation shows that the complete $v$-ary tree of height $h$ has leaves $L$ and total vertices/nodes $V$ given by
$L=v^{h}\mbox{ and }V=\frac{v^{h+1}-1}{h-1}.$ (14)
Taking the formulation of $L$ given in Eq. (14), to be the least integer upper bound of the right-hand side of Eq. (13), which is itself an upper-bound for $N_{\delta}$, and solving for $h$ yields:
$h=$ $\Big{\lceil}\log_{v}(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\Big{\rceil},$ (15)
where the integer ceiling was applied since $h$ must be an integer.
Let $\mathcal{L}$ be any set $v^{h}$ points in ${\mathcal{K}}$ containing the set $\{x_{i}\}_{i=1}^{N_{\delta}}$. Let $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ be any complete binary tree with leaves $\mathcal{L}$; note that, $\mathcal{L}\subseteq V$. By construction, and the computation in Eq. (14), $\mathcal{T}$ has $v^{h}$ leaves and $\frac{Lv-1}{h-1}$ nodes. ∎
For completeness, we include a minor modification of the proof of Proposition (Acciaio et al., 2023, Proposition 3.10), which allows for the approximation of uniformly continuous functions of arbitrarily low regularity. The original formulation of that result only allows for $\alpha$-Hölder function.
Proof of Lemma A.3.
If $f(x)=c$ for some constant $c>0$, then the statement holds with the neural network $\hat{f}(x)=c$, which can be represented as in (4) with $[d]=(n,m)$, where $A^{j}$ is the 00 matrix for all $j$, and the “$c$” in (4) is taken to be this constant $c$. Therefore, we henceforth only need to consider the case where $f$ is not constant. Let us observe that, if we pick some $x^{\star}\in\mathcal{K}$, then for any multi-index $[d]$ and any neural network $\hat{f}_{\theta}\in\mathcal{NN}_{[d]}^{\sigma}$, $\hat{f}_{\theta}(x)-f(x^{\star})\in\mathcal{NN}_{[d]}^{\sigma}$, since $\mathcal{NN}_{[d]}^{\sigma}$ is invariant to post-composition by affine functions. Thus, we represent $\hat{f}_{\theta}(x)-f(x^{\star})=\hat{f}_{\theta^{\star}}(x)$, for some $\theta^{\star}\in\mathbb{R}^{P([d])}$. Consequently:
$\sup_{x\in{\mathcal{K}}}\left|\|(f(x)-f(x^{\star}))-\hat{f}_{\theta^{\star}}(x)\|-\|f(x)-\hat{f}_{\theta}(x)\|\right|=0.$
Therefore, without loss of generality, we assume that $f(x^{*})=0$ for some $x^{*}\in{\mathcal{K}}$. By Benyamini & Lindenstrauss (2000, Theorem 1.12), there exists an $\omega$-uniformly continuous map $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ extending $f$.
Step 1 – Normalizing $\tilde{f}$ to the Unit Cube: First, we identify a hypercube “nestling” ${\mathcal{K}}$. To this end, let
$r_{\mathcal{K}}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{diam}({\mathcal{K}})\sqrt{\frac{n}{2(n+1)}}.$ (16)
By Jung’s Theorem (see (Jung, 1901)), there exists $x_{0}\in\mathbb{R}^{n}$ such that the closed Euclidean ball $\overline{\operatorname{Ball}_{(\mathbb{R}^{n},d_{n})}\left(x_{0},r_{\mathcal{K}}\right)}$ contains ${\mathcal{K}}$. Therefore, by Hölder’s inequality, we have that the $n$-dimensional hypercube $[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]$   [^3] For $x,y\in\mathbb{R}^{n}$ we denote by $[x,y]$ the hypercube defined by $\prod_{i=1}^{n}[x_{i},y_{i}]$.contains $\overline{B_{({\mathbb{R}^{n}},d_{n})}\left(x_{0},r_{\mathcal{K}}\right)}$, where $\bar{1}=(1,\dots,1)\in\mathbb{R}^{n}$. Consequently, ${\mathcal{K}}\subseteq[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]$. Let $\tilde{f}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}F|_{[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]}$, then $\tilde{f}\in C([x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}],\mathbb{R}^{m})$ is an $\omega$-uniformly continuous extension of $f$ to $[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]$.
Since ${\mathcal{K}}$ has at least two distinct points, then $r_{\mathcal{K}}>0$. Hence, the affine function
$T:\mathbb{R}^{n}\ni x\mapsto(2r_{\mathcal{K}})^{-1}(x-x_{0}+r_{\mathcal{K}}\bar{1})\in\mathbb{R}^{n}$
is well-defined, invertible, not identically 00, and maps $[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}-r_{\mathcal{K}}\bar{1}]$ to $[0,1]^{n}$. A direct computation shows that $g\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{f}\circ T^{-1}$ is also uniformly continuous, whose modulus of continuity $\tilde{\omega}:[0,\infty)\to[0,\infty)$ is given by
$\tilde{\omega}(t)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\omega(2r_{\mathcal{K}}\,t)$ (17)
for all $t\in[0,\infty)$. Furthermore, since for each $i=1,\dots,m$, define $\operatorname{pj}_{i}:\mathbb{R}^{m}\ni y\mapsto y_{i}\in\mathbb{R}$. Since each $\operatorname{pj}_{i}$ is $1$-Lipschitz then, for each $i=1,\dots,m$, the map $g_{i}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{pj}_{i}\circ g:[0,1]^{n}\to\mathbb{R}$ is also $\tilde{\omega}$-uniformly continuous. By orthogonality, we also note that $g(x)=\sum_{i=1}^{m}\,g_{i}(x)\,e_{i}$, for each $x\in\mathbb{R}^{n}$, where $e_{1},\dots,e_{m}$ is the standard orthonormal basis of $\mathbb{R}^{m}$; i.e. the $i^{th}$ coordinate of $e_{j}$ is $1$ if and only if $i=j$ and 00 is otherwise.
Step 2 – Constructing the Approximator: For $i=1,\dots,m$, let $\hat{f}_{\theta^{(i)}}\in\mathcal{NN}_{[d^{(i)}]}^{\sigma}$ for some multi-index $[d^{(i)}]=(d_{0}^{(i)},\dots,d_{J}^{(i)})$ with $n$-dimensional input layer and $1$-dimensional output layer, i.e. $d_{0}^{(i)}=n$ and $d_{J}^{(i)}=1$, and let $\theta^{(i)}\in\mathbb{R}^{P([d^{(i)}])}$ be the parameters defining $\hat{f}_{\theta^{(i)}}$. Since the pre-composition by affine functions and the post-composition by linear functions of neural networks in $\mathcal{NN}_{[d^{(i)}]}^{\sigma}$ are again neural networks in $\mathcal{NN}_{[d^{(i)}]}^{\sigma}$, we have that $g_{\theta^{(i)}}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\hat{f}_{\theta^{(i)}}\circ T^{-1}$ belongs to $\mathcal{NN}_{[d^{(i)}]}^{\sigma}$. Denote the standard basis of $\mathbb{R}^{m}$ by $\{e_{i}\}_{i=1}^{m}$. We compute:
$\sup_{x\in{\mathcal{K}}}\,\left\|f(x)-\sum_{i=1}^{m}\hat{f}_{\theta^{(i)}}(x)e_{i}\right\|$
$=$ $\sup_{x\in{\mathcal{K}}}\,\left\|\tilde{f}(x)-\sum_{i=1}^{m}\hat{f}_{\theta^{(i)}}(x)e_{i}\right\|$
$\leq$ $\sup_{x\in[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]}\,\left\|\tilde{f}(x)-\sum_{i=1}^{m}\hat{f}_{\theta^{(i)}}(x)e_{i}\right\|$
$=$ $\sup_{x\in[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]}\,\left\|\tilde{f}\circ T^{-1}\circ T(x)-\sum_{i=1}^{m}\hat{f}_{\theta^{(i)}}\circ T^{-1}\circ T(x)e_{i}\right\|$
$=$ $\sup_{u\in[0,1]^{n}}\,\left\|\sum_{i=1}^{m}g_{i}(u)\,e_{i}-\sum_{i=1}^{m}g_{\theta^{(i)}}(u)e_{i}\right\|$
$\leq$ $\sqrt{m}\max_{u\in[0,1]^{n}}\,\max_{1\leq i\leq m}\left|g_{i}(u)-g_{\theta^{(i)}}(u)\right|.$ (18)
Fix $\tilde{\varepsilon}>0$, to be determined below. For each $i=1,\dots,m$, depending on which assumption $\sigma$ satisfies, (Shen et al., 2022b, Theorem 1.1) (resp. (Shen et al., 2022a, Theorem 1) if $\sigma$ is as in Definition 10) imply that there is a neural network with activation function $\sigma^{\star}:\mathbb{R}\to\mathbb{R}$ satisfying
$\max_{u\in[0,1]^{n}}\,\left|g_{i}(u)-g_{\theta^{(i)}}(u)\right|<\tilde{\varepsilon}.$ (19)
Furthermore, the depth and width of these MLPs can be bounded above on a case-by-case basis as follows:
  1. (i)
    If $\sigma$ satisfies Definition A.1 then, setting each $\alpha=0$ implies that $\sigma_{0}(x)=\sigma^{\star}(x)$, as defined in Eq. (10); thus
    $J^{(i)}\leq 11\,\mbox{ and }\,\max_{1\leq j\leq J^{(i)}}d_{j}\leq 36n(2n+1)$
    In this case, we set $\tilde{\varepsilon}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\varepsilon/\sqrt{m}$; we have used Shen et al. (2022a, Theorem 1).
  2. (ii)
    If $\sigma$ satisfies Definition 3.1 then, setting each $\alpha=1$ implies that $\sigma_{0}(x)=\operatorname{ReLU}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\max\{0,x\}$; yielding
    $J^{(i)}\leq 18+2n+11\Big{\lceil}\Big{(}\frac{2r_{\mathcal{K}}}{\omega^{-1}\big{(}\varepsilon/(131\sqrt{n\,m})\big{)}}\Big{)}^{n/2}\Big{\rceil}\,\mbox{ and }\,\max_{1\leq j\leq J^{(i)}}d_{j}\leq 16\,\max\{n,3\}$
    in this case, we have employed Shen et al. (2022b, Theorem 1.1).
In either case, the estimate in Eq. (18) yields
$\max_{x\in{\mathcal{K}}}\,\left\|f(x)-\sum_{i=1}^{m}\hat{f}_{\theta^{(i)}}(x)e_{i}\right\|<\varepsilon.$
Step 3 – Assembling into an MLP:
Let $g_{1}\bullet g_{2}$ denote the component-wise composition of a univariate function $g_{1}$ with a multivariate function $g_{2}$.
If the activation function $\sigma$ is either in Definitions A.1 or 3.1, then it trivially implements the identity $I_{\mathbb{R}}$ on $\mathbb{R}$ by setting $\alpha=0$; i.e. $\sigma_{1}=I_{\mathbb{R}}$. Consequentially, for any $k\in\mathbb{N}_{+}$, if $I_{k}$ denotes the $k\times k$-identity matrix, then $I_{k}\sigma_{1}\bullet I_{k}\in\mathcal{NN}_{[d_{k}]}^{\sigma}$ with $P([d])=2k$, and $I_{k}\sigma_{1}\bullet I_{k}=1_{\mathbb{R}^{k}}$. Therefore, mutatis mutandis, $\mathcal{NN}_{[\cdot]}^{\sigma}$ satisfies the $c$-identity requirement with [^4] Formally, it satisfies what is the $1$-identity requirement, thus it satisfies the $c$-identity requirement for all integers $c\geq 2$. However, the authors of Cheridito et al. (2021) do not explicitly consider the extremal case where $c=1$. $c=2$, as defined in Cheridito et al. (2021, Definition 4). From there, mutatis mutandis, we may apply Cheridito et al. (2021, Proposition 5). Thus, there is a multi-index $[d]=(d_{0},\dots,d_{J})$ with $d_{0}=n$ and $d_{J}=m$, and a network $\hat{f}_{\theta}\in\mathcal{NN}_{[d]}^{\sigma}$ implementing $\sum_{i=1}^{m}\hat{f}_{\theta^{(i)}}e_{i}$, i.e.
$\sum_{i=1}^{m}\hat{f}_{\theta^{(i)}}e_{i}=\hat{f}_{\theta},$
such that $\hat{f}_{\theta}$’s depth and width are bounded-above, on a case-by-case basis, by
  1. (i)
    If $\sigma$ satisfies Definition A.1 then, setting each $\alpha=0$
    $J\leq$ $15\,m$
[…]
    In this case, we set $\tilde{\varepsilon}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\varepsilon/\sqrt{m}$.
  2. (ii)
    If $\sigma$ satisfies Definition 3.1 then, setting each $\alpha=0$ yields
    $J\leq$ $m\left(19+2n+11\Big{\lceil}\Big{(}\frac{2r_{\mathcal{K}}}{\omega^{-1}\big{(}\varepsilon/(131\sqrt{n\,m})\big{)}}\Big{)}^{n/2}\Big{\rceil}\right)$
[…]
Incorporating the definition of $r_{\mathcal{K}}$ in Eq. (16) and employing the inequality $\operatorname{diam}({\mathcal{K}})\leq 2\operatorname{rad}({\mathcal{K}})$ completes the proof. ∎
Proof of Lemma A.4.
Consider the tree $\tilde{\mathcal{T}}$ given by Lemma A.2. For each leaf $v_{i}$ of $\tilde{\mathcal{T}}$, we apply Lemma A.3 to deduce the existence of an MLP $\hat{f}_{\theta_{i}}$ with explicit depth and width estimates given by that lemma, satisfying the uniform estimate
$\max_{\|x-v_{i}\|\leq\delta}\,\|f(x)-\hat{f}_{\theta_{i}}(x)\|<\varepsilon.$ (20)
Let $\mathcal{T}$ be the same tree as $\tilde{\mathcal{T}}$ with leaves identified with $\{(v_{i},\theta_{i})\}_{i=1}^{L}$. ∎
We now prove our main technical version, namely Theorem A.5, which directly implies the special case recorded in Theorem 3.4.
Proof of Theorem A.5 (and this Theorem 3.4).
Applying Lemma A.4 with $\delta$ given as the solution of
$\Big{(}\frac{\delta 2^{3/2}n^{1/2}}{(n+1)^{1/2}\omega^{-1}(\varepsilon/131\,(nm)^{1/2})}\Big{)}^{n/2}\leq\Big{(}\frac{\delta 2^{3/2}n^{1/2}}{(2n^{1/2}\omega^{-1}(\varepsilon/131\,(nm)^{1/2})}\Big{)}^{n/2}=\varepsilon^{-r}.$ (21)
Solving Eq. (21) for $\delta$ implies that it is given by
$\delta=\frac{\varepsilon^{-2r/n}}{2}\,\omega^{-1}\biggl{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\biggr{)}.$
This completes the proof. ∎
Proof of Theorem 3.4.
Setting ${\mathcal{K}}=B_{n}(0,1)$, $r=1/2$, $\omega(t)=Lt$, and thus $\omega^{-1}(t)=L^{-1}\,t^{1/\alpha}$, in Theorem A.5 yields the conclusion. Finally, by the computation in Example 1, we have that $C\leq 2^{n+1}$; thus, $\log(C)=(n+1)\log(2)\leq 2n$. Noting that $c=1/\log(2)=1$ completes the proof. ∎
Remark B.1.
[…]
B.1 Proofs of Results Concerning VC dimension
We use the following lemma and its proof due to Gabriel Conant.
Lemma B.2 ((Conant, )).
Fix $n,L,d\in\mathbb{N}_{+}$. Let $\mathcal{H}$ be a non-empty set of functions from $\mathbb{R}^{n}$ to $\{0,1\}$ of VC dimension at-most $d$. Let $\mathcal{C}_{L}$ be the set of all ordered partitions (Voronoi diagrams) $(C_{l})_{l=1}^{\tilde{L}}$ covering $\mathbb{R}^{n}$, where $\tilde{L}\leq L$, and for which there exist distinct $p_{1},\dots,p_{\tilde{L}}\in\mathbb{R}^{n}$ such that: for each $l=1,\dots,\tilde{L}$
$C_{l}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{C}_{l}\setminus\bigcup_{s<l}\,\tilde{C_{s}}$ (22)
$\tilde{C}_{l}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{x\in\mathbb{R}^{n}:\,\|x-p_{l}\|=\min_{s=1,\dots,\tilde{L}}\,\|x-p_{s}\|\}.$
Let $\mathcal{H}_{L}$ be the set of functions from $\mathbb{R}^{n}$ to $\{0,1\}$ of the form
$f=\sum_{l=1}^{\tilde{L}}\,f_{l}\,I_{C_{l}}$
where $\tilde{L}\in\mathbb{N}_{+}$ with $\tilde{L}\leq L$, $f_{1},\dots,f_{\tilde{L}}\in\mathcal{H}$, and $(C_{l})_{l=1}^{\tilde{L}}\in\mathcal{C}_{L}$. Then, the VC dimension of $\mathcal{H}_{L}$ satisfies
$\operatorname{VC}(\mathcal{H}_{L})\leq 8L\log(\max\{2,L\})^{2}\,\big{(}\max\{d,2(n+1)\,(L-1)\,\log(3L-3)\}\big{)}.$
Proof of Lemma B.2.
Let us first fix our notation. Each $x_{0}\in\mathbb{R}^{n}\setminus\{0\}$ and $t\in\mathbb{R}$ defines a halfspace in $\mathbb{R}^{n}$ given by $HS_{x_{0},t}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{x\in\mathbb{R}^{n}:\,\langle x_{0},x\rangle\leq t\}$ (see (Boyd & Vandenberghe, 2004, Section 2.2.1) for details). We denote set of all halfspaces in $\mathbb{R}^{n}$ by $\mathcal{HS}_{n}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{HS_{x_{0},t}:\,\exists x_{0}\in\mathbb{R}^{n}\setminus\{0\}\,\exists t\in\mathbb{R}\}$. Consider the set $\mathcal{C}(L)$ of all $C\subseteq\mathbb{R}^{n}$ of the form
$C=\bigcap_{l=1}^{\tilde{L}-1}\,X_{l}$ (23)
for some positive integer $2\leq\tilde{L}\leq\max\{2,L\}$ and $X_{1},\dots,X_{\tilde{L}-1}\in\mathcal{HS}_{n}$.
Step 1 - Reformulation as Set of Sets
By definition of the powerset $2^{\mathbb{R}^{n}}$ of the set $\mathbb{R}^{n}$, each subset $A\subseteq\mathbb{R}^{n}$ can be identifies with a function (classifier) from $\mathbb{R}^{n}$ to $\{0,1\}$ via the bijection mapping any $X\in 2^{\mathbb{R}^{n}}$ to the binary classifier $I_{X}$ (i.e. the indicator function of the set $X$). Using this bijection, we henceforth identify both $\mathcal{H}$ and $\mathcal{H}_{L}$ with subsets of the powerset $2^{\mathbb{R}^{n}}$.
Under this identification, the class $\mathcal{H}_{L}$ can be represented as the collection of subsets $X$ of $\mathbb{R}^{n}$ of the form
$X=\bigcup_{l=1}^{\tilde{L}}\,H_{l}\cap C_{l},$ (24)
where $\tilde{L}\in\mathbb{N}_{+}$ satisfies $\tilde{L}\leq L$, and for each $l=1,\dots,\tilde{L}$ we have $H_{l}\in\mathcal{H}$ and $(C_{l})_{l=1}^{\tilde{L}}\in\mathcal{C}_{L}$ is of the form (22) for some distinct points $p_{1},\dots,p_{\tilde{L}}\in\mathbb{R}^{n}$.
Step 2 - VC Dimension of Voronoi Diagrams with at-most $L$ Cells
An element of $(C_{l})_{l=1}^{\tilde{L}}$ of $\mathcal{C}_{L}$ is, by definition, a Voronoi diagram in $\mathbb{R}^{n}$ and thus, Boyd & Vandenberghe (2004, Exercise 2.9) implies that each $C_{1},\dots,C_{\tilde{L}}$ is the intersection of $\tilde{L}-1\leq L-1$ halfspaces; i.e. $C_{1},\dots,C_{\tilde{L}}\in\mathcal{C}(L)$ (see Eq. (23)). Since $\mathcal{C}_{L}=\{\cap_{l=1}^{\tilde{L}}\,H_{i}:\,\exists\tilde{L}\in\mathbb{N}_{+}\,H_{1},\dots,H_{\tilde{L}}\in\mathcal{HS}_{n}\,\tilde{L}\leq L\}$ then Blumer et al. (1989, Lemma 3.2.3) implies that
$\operatorname{VC}(\mathcal{C}_{L})\leq$ $2\operatorname{VC}(\mathcal{HS}_{n})\,(L-1)\,\log(3L-3)$ (25)
$\leq$ $2(n+1)\,(L-1)\,\log(3L-3);$
the second inequality in Eq. (25) holds since $\operatorname{VC}(\mathcal{HS}_{n})=n+1$ by Shalev-Shwartz & Ben-David (2014, Theorem 9.3).
Step 3 - VC Dimension of The Class $\mathcal{H}_{L}$
Define $\mathcal{H}\cap\mathcal{C}(L)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{H\cap C:\,H\in\mathcal{H}\mbox{ and }C\in\mathcal{C}(L)\}$. Again using Blumer et al. (1989, Lemma 3.2.3), we have
[…]
$\leq$ $4\log(6)\,\big{(}\max\{\operatorname{VC}(\mathcal{H}),\operatorname{VC}(\mathcal{C}(L))\}\big{)}$
$\leq$ $4\log(6)\,\big{(}\max\{d,2(n+1)\,(L-1)\,\log(3L-3)\}\big{)}.$
Consider the set ~=$\tilde{\mathcal{H}}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{\cup_{l=1}^{\tilde{L}}\,H_{l}:\,\tilde{L}\in\mathbb{N}_{+},\,\tilde{L}\leq L,\,\forall l=1,\dots,\tilde{L},\,H_{l}\in\mathcal{H}\cap\mathcal{C}(L)\}$. Applying Blumer et al. (1989, Lemma 3.2.3), one final time yields
$\operatorname{VC}\big{(}\tilde{\mathcal{H}}\big{)}\leq$ $2\operatorname{VC}\big{(}\mathcal{H}\cap\mathcal{C}(L)\big{)}\,L\log(3L)$ (27)
$\leq$ $2\Big{(}4\log(6)\,\big{(}\max\{d,2(n+1)\,(L-1)\,\log(3L-3)\}\big{)}\Big{)}\,L\log(3L)$ (28)
$\leq$ $8L\log(\max\{2,L\})^{2}\,\big{(}\max\{d,2(n+1)\,(L-1)\,\log(3L-3)\}\big{)}$ (29)
where Eq. (28) held by the estimate in Eq. (26). Finally, since $\operatorname{VC}(A)\leq\operatorname{VC}(B)$ whenever $A\subseteq B$ for any set $B$ then since $\mathcal{H}_{L}\subseteq\tilde{\mathcal{H}}$ then Eq. (27)-(29) yields the desired conclusion. ∎
We may now derive Proposition 3.6 by merging Lemma B.2 and one of the main results of Bartlett et al. (2019).
Proof of Proposition 3.6.
Let $n,J,W,L\in\mathbb{N}_{+}$ and consider the (non-empty) set of real-valued functions $\mathcal{NN}_{J,W:n,1}^{\operatorname{PReLU}}$. By definition of the VC dimension of a set of real-valued functions, given circa Eq. (6), we have
$\operatorname{VC}\big{(}\mathcal{NN}_{J,W:n,1}^{\operatorname{PReLU}}\big{)}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}$ $\operatorname{VC}\big{(}I_{(0,\infty)}\circ\mathcal{NN}_{J,W:n,1}^{\operatorname{PReLU}}\big{)}$
$\operatorname{VC}\big{(}\mathcal{NP}_{J,W,L:n,1}^{\operatorname{PReLU}}\big{)}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}$ $\operatorname{VC}\big{(}I_{(0,\infty)}\circ\mathcal{NP}_{J,W,L:n,1}^{\operatorname{PReLU}}\big{)}.$ (30)
By Bartlett et al. (2019, Theorem 7), we have that
$\operatorname{VC}(I_{(0,\infty)}\circ\mathcal{NN}_{J,W:n,1}^{\operatorname{PReLU}})\leq D^{\star}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{\lceil}J+(J+1)\,W^{2}\,\log_{2}\big{(}e\,4(J+1)\,W\log_{2}(e2(J+1)W)\big{)}\big{\rceil}.$ (31)
Therefore, applying Lemma B.2 with $\mathcal{H}=\big{(}I_{(0,\infty)}\circ\mathcal{NN}_{J,W:n,1}^{\operatorname{PReLU}}\big{)}$ yields the estimate
$\operatorname{VC}\big{(}I_{(0,\infty)}\circ\mathcal{NP}_{J,W,L:n,1}^{\operatorname{PReLU}}\big{)}\leq 8L\log(\max\{2,L\})^{2}\,\big{(}\max\{D^{\star},2(n+1)\,(L-1)\,\log(3L-3)\}\big{)}$ (32)
Combining Eq. (32) and the definition (30) yields the bound. In particular,
$\operatorname{VC}\big{(}\mathcal{NP}_{J,W,L:n,1}^{\operatorname{PReLU}}\big{)}\in\mathcal{O}\big{(}L\log(L)^{2}\,\max\{nL\log(L),JW^{2}\log(JW)\}\big{)}$
yielding the second conclusion. ∎
Proof of Proposition 3.7.
We argue by contradiction. Suppose that $\mathcal{F}$ has finite VC dimension $\operatorname{VC}(\mathcal{F})$. Then, Shen et al. (2022b, Theorem 2.4) implies that there exists a $1$-Lipschitz map $f:[0,1]^{n}\to\mathbb{R}$ such that does not exist a strictly positive $\varepsilon\in(0,\operatorname{VC}(\mathcal{F})^{-1/n}/9)$ satisfying such that
$\inf_{\hat{f}\in\mathcal{F}}\,\sup_{x\in[0,1]^{n}}\,|\hat{f}(x)-f(x)|\leq\varepsilon.$ (33)
However, Shen et al. (2022a, Theorem 1) implies that, for every $1$-Lipschitz function, in particular for $f$, and for each $\tilde{\varepsilon}>0$ there exists a $\hat{f}_{\tilde{\varepsilon}}\in\mathcal{F}$ satisfying
$\sup_{x\in[0,1]^{n}}\,|\hat{f}_{\tilde{\varepsilon}}(x)-f(x)|\leq\tilde{\varepsilon}.$ (34)
Setting $\tilde{\varepsilon}=\operatorname{VC}(\mathcal{F})^{-1/n}/18$ yields a contradiction as Eq. (33) and Eq. (34) cannot both be simultaneously true. Therefore, $\mathcal{F}$ has infinite VC dimension. ∎
Appendix C The Curse of Irregularity
We now explain why learning Hölder functions of low regularity ($(1,1/d)$-Hölder) functions on the real line segment $[0,1]$ is equally challenging as learning regular functions ($1$-Lipschitz) on $[0,1]^{n}$.
C.1 Hölder Functions
Fix $n,m\in\mathbb{N}$ and let $\mathcal{X}\subset\mathbb{R}^{n}$ be non-empty and compact of diameter $D$. Fix $0<\alpha\leq 1$, $L\geq 0$, and let $f:\mathcal{X}\rightarrow\mathbb{R}^{m}$ be $(\alpha,L)$-Hölder continuous, meaning
$\|f(x)-f(y)\|\leq\,L\|x-y\|^{\alpha}$
holds for each $x,y\in\mathcal{X}$. For any $L\geq 0$ and $0<\alpha\leq 1$, we denote set of all $(\alpha,L)$-Hölder functions from $\mathcal{X}$ to $\mathbb{R}^{n}$ is denoted by $C^{\alpha}([0,1]^{n},\mathbb{R};L)$.
We focus on the class of locally Hölder functions since they are generic, i.e. universal, amongst all continuous functions by the Stone–Weierstrass theorem. In this case, Hölder functions are sufficiently rich to paint a full picture of the hardness to approximate arbitrary Hölder functions either by MLPs against the proposed model.
In contrast to smaller generic function classes, such as polynomials, Hölder functions provide more freedom in experimentally visualizing our theoretical results. This degree of freedom is the parameter $\alpha$, which modulates their regularity. As $\alpha$ tends to 00 the Hölder functions become complicated and when $\alpha=1$ the Rademacher-Stephanov theorem, see (Federer, 1978, Theorems 3.1.6 and 3.1.9) characterizes Lebesgue almost-everywhere differentiable functions as locally $(1,L)$-Hölder maps. Note that $(1,L)$-Hölder functions are also called $L$-Lipschitz maps and, in this case, $L=\sup_{x}\,\|\nabla f(x)\|_{op}$ where the supremum is taken over all points where $f$ is differentiable and where $\|\cdot\|_{op}$ is the operator norm.
C.2 The Curse of Irregularity
The effect of low Hölder regularity, i.e. when $\alpha\approx 0$, has the same effect as high-dimensionality on the approximability of arbitrary $\alpha$-Hölder functions. This is because any real-valued model/hypothesis class $\mathcal{F}_{1}$ of functions on $\mathbb{R}$ approximating an arbitrary $(\frac{1}{d},1)$-Hölder functions By (Shen et al., 2022b, Theorem 2.4), we have the lower minimax bound: if for each $\varepsilon>0$ we have “the curse of irregularity”
[…]
where the supremum is taken over all $f\in C^{1/d}([0,1],\mathbb{R};1)$. The familiar curse of dimensionality also expresses the hardness to approximate an arbitrary $1$-Lipschitz ($(1,1)$-Hölder), thus relatively regular, function on $[0,1]^{n}$. As above, consider any model/hypothesis class $\mathcal{F}_{2}$ of real-valued maps on $\mathbb{R}^{d}$ then, again using (Shen et al., 2022b, Theorem 2.4), one has the lower-bound
[…]
where the supremum is taken over all $f\in C^{1}([0,1]^{n},\mathbb{R};1)$. Comparing Eq. (35) and Eq. (36), we find that the difficulty of uniformly approximating an arbitrary low-regularity ($(\frac{1}{d},1)$-Hölder) function on a $1$-dimensional domain is roughly just as complicated as approximating a relatively regular ($1$-Lipschitz) function on a high-dimensional domain.
Incorporating these lower bounds with the lower-bound in Eq. (7), we infer that the minimum number layers $(L)$ and minimal width $(W)$ of each MLP approximating a low-regularity function on the low-dimensional domain $[0,1]$ is roughly the same as the minimal number of layers and width of an MLP approximating a high-regularity map on the high-dimensional domain $[0,1]^{n}$.
Appendix D Experimental details
We include here experimental details, we refer to the source code in the supplementary material for more details.
D.1 Formulations of the Ackley and Rastrigin functions
Let us note $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$ the $n$-dimensional representation of a sample, we use the following formulation of the Ackley function:
$\operatorname{Ackley}(x)=20+\exp(1)-a\exp\left(-b\sqrt{\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}}\right)-\exp\left(\frac{1}{n}\sum_{i=1}^{n}\cos(2\pi x_{i})\right)$ (37)
where $a=20$ and $b=0.2$. We also use the following formulation of the Rastrigin function:
$\operatorname{Rastrigin}(x)=\sum_{i=1}^{n}x_{i}^{2}+10\left(n-\sum_{i=1}^{n}\cos(2\pi x_{i})\right)$ (38)
D.2 Architectures and hyperparameters
D.2.1 Neural pathways
We consider that the width of our neural pathways is $w=1000$. In other words, each hidden layer of our neural pathways contains a linear matrix of size $w\times w$.
In the regression task, our neural pathways contain 3 hidden layers and we use a learnable PReLU as activation function. We use Adam (Kingma & Ba, 2014) as optimizer with a learning rate of $410^{-}$ 4 and other default parameters of Pytorch.
In the classification task, we follow the setup of Oquab et al. (2023) and then use AdamW (Loshchilov & Hutter, 2019) as optimizer with a learning rate of $10^{-3}$ and other default parameters of Pytorch, our neural pathways contain 4 hidden layers in the classification task, and we apply BatchNorm1d before the PReLU activation function.
D.2.2 Baseline
The baseline has the same architecture as the neural pathways above. However, if we note $K$ the number of prototypes and assume that $\sqrt{K}$ is a natural number, the width of the baseline is $w\sqrt{K}$ so that the number of hidden parameters is the same as all the neural pathways combined.
In our experiments, we set $K=4$, so $\sqrt{K}=2$.

### v2

@@
-Breaking the Curse of Dimensionality with Distributed Neural Computation
+Approximation Rates and VC-Dimension Bounds for (P)ReLU MLP Mixture of Experts
 Abstract
-We present a theoretical approach to overcome the curse of dimensionality using a neural computation algorithm which can be distributed across several machines.
-Our modular distributed deep learning paradigm, termed neural pathways, can achieve arbitrary accuracy while only loading a small number of parameters into GPU VRAM.
-Formally, we prove that for every error level $\varepsilon>0$ and every Lipschitz function $f:[0,1]^{n}\to\mathbb{R}$, one can construct a neural pathways model which uniformly approximates $f$ to $\varepsilon$ accuracy over $[0,1]^{n}$ while only requiring networks of $\mathcal{O}(\varepsilon^{-1})$ parameters to be loaded in memory and $\mathcal{O}(\varepsilon^{-1}\log(\varepsilon^{-1}))$ to be loaded during the forward pass.
-This improves the optimal bounds for traditional non-distributed deep learning models, namely ReLU MLPs, which need $\mathcal{O}(\varepsilon^{-n/2})$ parameters to achieve the same accuracy.
-The only other available deep learning model that breaks the curse of dimensionality is MLPs with super-expressive activation functions.
-However, we demonstrate that these models have an infinite VC dimension, even with bounded depth and width restrictions, unlike the neural pathways model.
-This implies that only the latter generalizes.
-Our analysis is validated experimentally in both regression and classification tasks, demonstrating that our model exhibits superior performance compared to larger centralized benchmarks.
-Distributed Computing, Approximation Theory, Deep Learning, VC Dimension.
+Mixture-of-Experts (MoEs) can scale up beyond traditional deep learning models by employing a routing strategy in which each input is processed by a single “expert” deep learning model.
+This strategy allows us to scale up the number of parameters defining the MoE while maintaining sparse activation, i.e., MoEs only load a small number of their total parameters into GPU VRAM for the forward pass depending on the input.
+In this paper, we provide an approximation and learning-theoretic analysis of mixtures of expert MLPs with (P)ReLU activation functions.
+We first prove that for every error level $\varepsilon>0$ and every Lipschitz function $f:[0,1]^{n}\to\mathbb{R}$, one can construct a MoMLP model (a Mixture-of-Experts comprising of (P)ReLU MLPs) which uniformly approximates $f$ to $\varepsilon$ accuracy over $[0,1]^{n}$, while only requiring networks of $\mathcal{O}(\varepsilon^{-1})$ parameters to be loaded in memory.
+Additionally, we show that MoMLPs can generalize since the entire MoMLP model has a (finite) VC dimension of $\tilde{O}(L\max\{nL,JW\})$, if there are $L$ experts and each expert has a depth and width of $J$ and $W$, respectively.
 1 Introduction
-The curse of dimensionality is a core obstruction in the foundations of machine learning.
-From the theoretical lens, given by the universal approximation theorem, the curse of dimensionality implies that every standard deep learning model approximately implementing an arbitrary almost-everywhere differentiable function on the $n$-dimensional cube $[0,1]^{n}$, up to an error of $\varepsilon>0$, requires at least $\Omega(\varepsilon^{-n/2})$ parameters (Yarotsky, 2018; Kratsios & Papon, 2022; Shen et al., 2021, 2022b).
-In practice, model upscaling is limited by the available device memory; Video RAM (VRAM) for GPUs.
-This happens either in training, during the backward pass when computing the Jacobian of the deep learning model, or even in the forward pass if the model is large enough.
-Currently, the curse of dimensionality is only known to be breakable in either one of two restrictive scenarios:
-Table 1: Overview of Main Results: Depth and width of the MLP architecture $\hat{f}_{\theta}$ and the number of active parameters in GPU VRAM (Params.
-in GPU) required to approximate an $\alpha$-Hölder function.
-The maximum number of parameters required during a forward pass, for any given input, are also recorded (Forward Pass).
-The optimal approximation rates for the standard ReLU MLP are derived in Yarotsky (2017); the optimality of which is expressed in terms of VC dimension in Shen et al.
+With the advent of large foundation models, the need to scale deep learning models beyond what is feasible to process on a single machine has become exceedingly relevant.
+Mixture of Experts (MoE) models present a solution to this problem via a sparse activation strategy.
+In MoEs, every input is first routed to one amongst a large number of expert deep learning models and then processed therewith.
+This allows MoEs to scale up, while maintaining a low/constant computational cost during the forward pass since only a subset of the overall model needs to be loaded into GPU video random-access memory (VRAM) for any given input.
+This has led to MoEs such as Mixtral (Jiang et al., 2024), Gemini (Google, 2023), and several others, e.g.
+(Jacobs et al., 1991; Jordan & Xu, 1995; Meila & Jordan, 2000; Shazeer et al., 2017; Guu et al., 2020; Lepikhin et al., 2021; Fedus et al., 2022; Barham et al., 2022; Majid & Tudisco, 2024), to become a viable solution in scaling up large language models (Radford et al., 2018; Brown et al., 2020).
+However, the analytical and statistical underpinnings of MoEs in deep learning are comparatively less understood in contrast to empirical investigations.
+This paper adds to the theoretical understanding of this subject by studying MoEs whose experts are (small) multilayer perceptrons (MLPs) with (P)ReLU activation function (MoMLPs).
+A key feature of MoEs is that they can maintain a small/fixed computational cost during the forward pass, for any given input $x\in[0,1]^{n}$, even if the overall model complexity may be large.
+Our main result (Theorem 4.1) analyzes the complexity of MoEs when uniformly approximating an arbitrary Lipschitz (Lebesgue) almost-everywhere continuously differentiable function $f:[0,1]^{n}\to\mathbb{R}^{m}$ by an MoMLP with (P)ReLU activation function to any prespecified error $\varepsilon>0$.
+We focus on the trade-off between the maximum number of parameters loaded into VRAM by any expert model $\{\hat{f}_{l}\}_{l=1}^{\ell}$ while predicting from any given input, against the total number of experts required to maintain that constant number of activated parameters in the forward pass.
+Summarized in Table 1, our main result shows that a constant active complexity in the forward pass can be maintained amongst all experts, but at the cost of an exponentially large number of locally-specialized experts $\{\hat{f}\}_{l=1}^{\ell}$ and regions of specialization $\{C_{l}\}_{l=1}^{\ell}$.
+Our complexity estimates are approximately optimal as they nearly match the Vapnik-Chervonenkis (VC) lower bounds derived in Shen et al.
+(2021).
+That is, the uniform approximation of an arbitrary such $f$ on $[0,1]^{n}$, with an an error of $\varepsilon>0$, requires at least $\Omega(\varepsilon^{-n/2})$ total model parameters (Yarotsky, 2018; Kratsios & Papon, 2022; Shen et al., 2021, 2022b).
+It is here where MoEs have an advantage since not all of these parameters need to be loaded into active memory for any given input; thus MoEs are genuinely sparsely activated.
+From the statistical learning perspective, a key property of MoEs (e.g.
+the top MoMLP model) is that they can maintain a given level of activation in the forward pass while the entire MoE model can maintain a finite VC-dimension (Theorem 4.2).
+This is key, for instance, in classification applications, as the fundamental theorem of PAC learning (see e.g.
+(Shalev-Shwartz & Ben-David, 2014, Theorem 6.7) or the results of Blumer et al.
+(1989); Hanneke (2016); Brukhim et al.
+(2022)) implies that such a machine learning model generalizes beyond the training data if and only if it has finite VC dimension.
+Summary of Contributions
+Table 1 summarizes our main contributions, both to the approximation theory and learning theory of MoE models, in the context of the toy mixture of (P)ReLU MLP models.
+All results illustrate the trade-off between individual (expert) complexity and the complexity shared across the set of experts when uniformly approximating a target $\alpha$-Hölder function, where $0<\alpha\leq 1$.
+Note that $\alpha=1$ is equivalent to the target function being differentiable (Lebesgue) almost-everywhere.
+Our approximation theoretic result (Theorem 4.1) records the number of parameters required to perform a uniform approximation on a high-dimensional Euclidean space on $[0,1]^{n}$.
+Our statistical learning theoretic result (Theorem 4.2) yields a bound on the VC dimension of the entire class of MoEs with just enough approximation power to perform this approximation.
+Table 1: Parametric Complexity and VC-Dimension of MoMLP model with no.
+experts-to-expert complexity parameter $r\in\mathbb{R}$; performing an $\varepsilon>0$ approximation of an $\alpha$-Hölder function $f:[0,1]^{n}\to\mathbb{R}$; for $n\in\mathbb{N}$.
+When $r\geq 0$ more model complexity is distributed across many “small experts”.
+When $r<0$ fewer experts define the MoE and, as a result, each expert MLP must depend on more parameters such that the entire MoE obtain an accurate approximation of the target function.
+Parameter Estimate
+No. Expert Parameters $\mathcal{O}(\max\{1,\varepsilon^{-r}\})$
+No. Experts $\mathcal{O}\big{(}\max\{1,\varepsilon^{2r/n-1/\alpha}\}\big{)}$
+Routing Complexity $\mathcal{O}\big{(}\max\{1,(\frac{2r}{n}-\frac{1}{\alpha})\log(\varepsilon)\}\big{)}$
+VC Dimension MoE $\tilde{\mathcal{O}}\big{(}\max\{1,\varepsilon^{2r/n-1/\alpha}\}\,\max\{\varepsilon^{2r/n-1/\alpha},\varepsilon^{-r}\}\big{)}$
+Observe that, setting $r=\frac{2}{n}$ in Table 1, yields for ReLU MLPs derived in Yarotsky (2017); the optimality of which is expressed in terms of VC dimension in Shen et al.
 (2022b).
-The rates for the non-Lipschitz continuous “unstable MLP model” are derived in Shen et al.
-(2022a).
-The VC dimension of the class of ReLU MLPs with bounded depth and width is computed in Bartlett et al.
-(2019), the (finite) VC dimension of our neural pathways model, with bounded depth, width, and number of parts, is shown in Proposition 3.6, and the infinite VC dimension of the class of unstable MLPs is shown in Proposition 3.7.
-Model Depth Width Params.
-in (V)RAM Forward Pass VC Dim Ours $\mathcal{O}(\varepsilon^{-1})$ $\mathcal{O}(1)$ ${\color[rgb]{0.09,0.45,0.27}\definecolor[named]{pgfstrokecolor}{rgb}{0.09,0.45,0.27}\pgfsys@color@rgb@stroke{0.09}{0.45}{0.27}\pgfsys@color@rgb@fill{0.09}{0.45}{0.27}{\mathcal{O}(\varepsilon^{-1})}}$ $\mathcal{O}\big{(}\frac{n}{\alpha}\log(1/\varepsilon)/\varepsilon\big{)}$ Finite ReLU MLP ${\color[rgb]{0.82,0.1,0.26}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.1,0.26}\pgfsys@color@rgb@stroke{0.82}{0.1}{0.26}\pgfsys@color@rgb@fill{0.82}{0.1}{0.26}{\mathcal{O}(\varepsilon^{-n/(2\alpha)})}}$ $\mathcal{O}(1)$ ${\color[rgb]{0.82,0.1,0.26}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.1,0.26}\pgfsys@color@rgb@stroke{0.82}{0.1}{0.26}\pgfsys@color@rgb@fill{0.82}{0.1}{0.26}{\mathcal{O}\big{(}\varepsilon^{-\frac{n}{2\alpha}}\big{)}}}$ $\mathcal{O}\big{(}\varepsilon^{-\frac{n}{2\alpha}}\big{)}$ Finite Unstable MLP $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Infinite
-Exceptionally Regular Target Functions.
-The curse of dimensionality can only be broken when approximating members of highly restrictive classes of target functions (Barron, 1993; Mhaskar et al., 2017; Suzuki, 2019; Opschoor et al., 2022; Galimberti et al., 2022; Gonon et al., 2023).
+Likewise, the VC dimension of the MoMLP is roughly equal to that of ReLU MLPs computed in Bartlett et al.
+(2019).
+2 Related Work
+Deep Learning Models with Few Parameters in Active Memory.
+Deep learning models with highly oscillatory “super-expressive” activation functions (Yarotsky & Zhevnerchuk, 2020; Yarotsky, 2021; Zhang et al., 2022) are known to achieve dimension-free approximation rates, thus effectively require a (relatively) feasible number of parameters to be loaded into VRAM during the forward pass.
+As we will see in Proposition 4.4, many of these models have an infinite VC dimension even when they are restricted to having a bounded depth and width; see Jiao et al.
+(2023, Lemma 3.1) for $\operatorname{ReLU-Sin-2^{x}}$-networks.
+Their unbounded VC dimension implies that the classifiers implemented by these models do not generalize on classification problems.
+Thus, the real performance of these models does not need to achieve the approximation-theoretic optima since they can only learn from a finite number of noisy training instances.
+Alternatively, a feasible number of parameters in deep learning models with standard activation functions may be guaranteed by restricting classes of well-behaved target functions such as Barron functions (Barron, 1993), functions of mixed-smoothness (Suzuki, 2019), highly smooth functions (Mhaskar, 1996; Galimberti et al., 2022; Gonon et al., 2023; Opschoor et al., 2022), convex functions (Bach, 2017), functions with compositional structure (Mhaskar et al., 2017), or other restricted classes.
 However, there are generally no verifiable guarantees that a target function encountered in practice has the necessary structure for these desired approximation theorems to hold.
-Generalization Problems for Super-expressive Multilayer Perceptions (MLPs).
-Deep learning models with highly oscillatory “super-expressive” activation functions (Yarotsky & Zhevnerchuk, 2020; Yarotsky, 2021; Zhang et al., 2022) boast approximation rates which are not cursed by dimensionality.
-However, as we will see in Proposition 3.7, many of these models have an infinite VC dimension even when they are restricted to having a bounded depth and width.
-Their unbounded VC dimension implies that the classifiers implemented by these models, by post-composition with $I_{(0,\infty)}$, do not generalize due to the classical results of Blumer et al.
-(1989); Hanneke (2016); Brukhim et al.
-(2022).
-Thus, the real performance of these models need not achieve the approximation-theoretic optima since they can learn from a finite number of training instances.
-This paper demonstrates that distributing neural computation offers a feasible third option for overcoming the curse of dimensionality in deep learning.
-Importantly, unlike the aforementioned cases, we show that distributed deep learning does not require high regularity of the target function being approximated, nor does it rely on models with unorthodox activation functions.
-Our deep learning algorithm, termed neural pathways, is distributed across the storage space (i.e., offline memory) of the machine.
-It only requires a small portion of the model to be loaded into VRAM at any given time, while the entire model is distributed over the machine’s (pragmatically infinite) storage space.
-In essence, each point in the input space is assigned to a trainable prototype (or landmark or representative point).
-Each of these prototypes is linked with a small neural network, specifically designed to approximate the target function for inputs in proximity to its associated prototype.
-During the forward pass, the input is initially directed to the appropriate prototype, after which the corresponding network is loaded into memory and employed to generate a prediction.
-Our algorithm follows the principle of route-first and predict-second via locally specialized neural networks.
-This approach mimics how stimuli activate only specialized neural pathways in the brain, without necessitating the involvement of the entire brain to process any single input.
-2 Neural Pathways
-In this section, we present the motivation behind the neural pathways framework proposed in this paper, and we describe it from an algorithmic perspective.
-The primary goal is to provide the reader with an intuitive understanding of the overall procedure before delving into a more involved mathematical analysis in Section 3, which is approached through the lens of approximation theory.
-Motivation.
-Our model formulation is loosely inspired by the cognitive sciences, and more specifically, by the fact that neuroscientists have observed that the brain does not activate completely to process each input, but rather uses specific regions of the overall network at a time, also known as brain functional specialization (Mahon & Cantlon, 2011).
-On the other hand, from a computational perspective, our motivation is to increase the size of deep learning models regardless of GPU memory capacity.
-We theoretically study the effect and feasibility of scaling up models by combining smaller distributed networks specialized to approximate specific regions of the input space.
-This differs from simply increasing the parameter count of a single large network, which is constrained by available VRAM.
-2.1 Model overview
-We can visualize our framework as a tree, with each leaf corresponding to a prototype and an associated neural network.
-In our case, we work with MLPs, but the network need not be a specific architecture.
-We refer to the combination of each prototype and its associated neural network as a neural pathway within our model.
-We refer to the overall algorithmic construction as neural pathways in plural.
-We can think of neural pathways as a framework for neural computation rather than a specific model since this approach could be applied to any other common neural networks such as CNNs (LeCun et al., 1989) or Transformers (Vaswani et al., 2017), for instance.
-Similarly, we propose specific algorithms to discover the prototypes for our neural pathways, but these could be substituted with enhanced approaches in the future.
-However, the main goal of this work is not to adapt the framework to a variety of specialized machine learning tasks but rather to provide its initial formulation and, particularly, to analyze its mathematical feasibility in terms of the curse of dimensionality.
-The neural pathways algorithm is divided into two stages: first, we discover the prototype locations, and then, we optimize the neural pathways associated with each prototype.
-In step one, our framework learns how to subdivide the input space, and in step two, we focus on neural network parameter optimization while keeping the input space partitioning fixed.
-The mathematical soundness of this approach is later presented as part of our main theoretical results in Section 3.
-A schematic of the algorithm workflow is depicted in Figure 1.
-Figure 1: $1$) The distance from each input $x$ to all prototypes $p_{1},\dots,p_{8}$ is queried.
-$2$) The network ($\hat{f}_{2}$ in the figure) assigned to the nearest prototype ($p_{2}$ in the figure) is loaded onto the GPU and used for prediction.
-2.2 Training Algorithm
-We now provide an explanation for the training algorithm.
-As discussed, we break down the algorithm into two parts: discovering the prototypes and training the networks.
-Conceptually, the prototypes define where in the input space the networks are located, or in other words, where in the input space we expect each of the networks to have the best performance.
-During inference, we will route a given input to the appropriate network based on its nearest prototype.
-In essence, each network learns to approximate a specific region of the overall input domain.
-Discovering prototypes.
-In principle, we may not know how to partition the input space.
-One approach is to utilize standard clustering algorithms like $K$-means, but this might be suboptimal for the downstream task unless we are already operating in a structured latent space, such as those found in pre-trained models (further discussion on this is available in Section 5).
-Another way is to learn it via gradient descent by optimizing the location of the prototypes for a specific task by following the gradient of the downstream loss.
-At the beginning of training, we have $\hat{F}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\hat{f}_{1}(x),\dots,\hat{f}_{K}(x))$ which contains a collection of $K$ randomly initialized shallow or small networks (i.e., much smaller than our neural pathways described later).
-In this first step, we assume that we are able to load all randomly initialized networks into our GPU memory.
-In particular, this is true because we use small networks with few parameters, which we will later “deepen” in the next step by adding additional hidden layers.
-We initialize the prototypes $p\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(p_{1},\dots,p_{K})$ randomly from a uniform distribution within the bounds of our training dataset input samples.
-We use the following expression to train the location of our prototypes $\{p_{k}\}_{k=1}^{K}$ in $\mathbb{R}^{n}$ by minimizing the energy:
-$\sum_{(x,y)\in\mathcal{D}}\,\ell\big{(}\operatorname{softmax}\big{(}-\|x-p_{i}\|_{i=1}^{K}\big{)}^{\top}\,\hat{F}(x)\,,y\big{)}.$ (1)
-where the loss $\ell$ is task-specific; for example, one could use mean squared error for regression and cross-entropy for classification.
-The $\operatorname{softmax}$ weights the importance of the prediction of each of the neural pathways in $\hat{F}$ for a given input $x$, as a function of the input’s distance to the prototypes, $\|x-p_{i}\|_{i=1}^{K}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\|x-p_{1}\|,...,\|x-p_{K}\|)$.
-Both the locations of the prototypes and the shallow randomly initialized neural networks assigned to them are optimized.
-Deepening the Networks.
-After the initial training phase, we enhance the networks by incorporating additional layers.
-Specifically, we introduce linear layers with weights initialized to the identity matrix and bias set to zero, just before the final output layer of each network.
-To encourage gradient flow in these new layers during the subsequent training stage, we slightly perturb this initialization with small Gaussian noise.
-This approach is driven by the fact that in the second training stage, each neural pathway can be optimized in a distributed manner.
-Consequently, we can work with larger networks without the need to load all of them simultaneously into our GPU, allowing for more model parameters.
-During the first stage, we have already optimized our networks alongside the prototype locations, converging towards a minimum.
-By initializing the networks with the new layers close to the identity, we can ensure that their output at the start of the second stage of training is similar to that produced by the original networks.
-This allows us to smoothly continue the optimization process from the point where we previously halted.
-Neural Pathway Training. Once prototype locations have been fixed we can independently train neural pathways $\hat{f}_{1},\dots,\hat{f}_{K}$ by minimizing for all $k\in\{1,\dots,K\}$:
-($\underset{\underset{k\in\operatorname*{arg\,min}_{j\in\{1,\dots,K\}}\{\|x-p_{j}\|\}}{(x,y)\in\mathcal{D}}}{\sum}\,\ell(\hat{f}_{k}(x),y)$ (2)
-over all the networks. We optimize the neural pathway network $\hat{f}_{k}$ for training data points that are closest to prototype $p_{k}$. The training procedure is summarized in Algorithm 1.
-Algorithm 1 Neural Pathways Training.
-Training data $\mathcal{D}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(x_{j},y_{j})\}_{j=1}^{N}$, no. of prototypes $K\in\mathbb{N}_{+}$, loss function $\ell$.
-Discovering Prototypes:
-$(\hat{F},p)\leftarrow{\operatorname*{arg\,min}_{\hat{F},p}}\,\displaystyle\sum_{(x,y)\in\mathcal{D}}\,\ell\left(\operatorname{softmax}\left(x|p\right)^{\top}\hat{F}(x),y\right)$
-Deepen networks:
-    For $k=1,\dots,K$:
-     $\hat{f}_{k}\leftarrow\textrm{deepen}(\hat{f}_{k})$
-Neural Pathway Training:
-   For $k=1,\dots,K$:
-     $\hat{f}_{k}\leftarrow\underset{\hat{f}_{k}}{\operatorname*{arg\,min}}\,\underset{\underset{k\in\operatorname*{arg\,min}_{j}\{\|x-p_{j}\|\}}{(x,y)\in\mathcal{D}}}{\sum}\,\ell(\hat{f}_{k}(x),y)$
-return Neural pathway parameters $\{\hat{f}_{k}\}_{k=1}^{K}$ and prototype locations $\{p_{k}\}_{k=1}^{K}$.
-Inference.
-At inference time, each test sample $x$ is assigned to its nearest prototype $p_{k}$ where $k\in\operatorname*{arg\,min}_{j\in\{1,\dots,K\}}\{\|x-p_{j}\|\}$ and the prediction is made by the $k$-th neural pathway $\hat{f}_{k}$.
-3 Mathematical Analysis of Neural Pathways
-We now analyze the approximation capacity of the neural pathways framework, when applied to MLPs with trainable PReLU activation functions.
-3.1 Preliminaries
-Before presenting our main result, Theorem 3.4, we first formalize the necessary mathematical terminology.
-Notation.
+Universal Approximation in Deep Learning.
+Several results have recently considered the expression power of deep learning models.
+These include universal approximation guarantees for MLPs (Cybenko, 1989; Hornik et al., 1989; Lu et al., 2017; Suzuki, 2019; Yarotsky, 2017, 2018; Voigtlaender & Petersen, 2019; Bolcskei et al., 2019; Gühring et al., 2020; De Ryck et al., 2021; DeVore et al., 2021; Daubechies et al., 2022; Kratsios & Zamanlooy, 2022; Zhang et al., 2022; Opschoor et al., 2022; Zamanlooy & Kratsios, 2022; Shen et al., 2022b; Cuchiero et al., 2023; Voigtlaender, 2023; Benth et al., 2023; Mao & Zhou, 2023; Yang & Zhou, 2024), CNNs (Petersen & Voigtlaender, 2020; Yarotsky, 2022), spiking neural networks (Neuman & Petersen, 2024), residual neural networks (Tabuada & Gharesifard, 2021), transformers (Yun et al., 2019, 2020; Kratsios & Papon, 2022; Fang et al., 2023), random neural networks (Gonon et al., 2023), recurrent neural network models (Grigoryeva & Ortega, 2018; Gonon & Ortega, 2021; Hutter et al., 2022; Galimberti et al., 2022; hoon Song et al., 2023), and several others.
+In each these cases, one typically considers the expressivity of a single “expert” model and not a mixture thereof.
+Our analysis can be customized to any of these settings to yield analogues of Theorem 4.1.
+Foundations of MoEs.
+MoE models have been heavily studied since their inception.
+Most results have focused on identifying the correct expert to best route any given input to (Teicher, 1960, 1963; Wang et al., 1996), the construction of effective routing mechanisms (Wang et al., 2017) selection (Wang et al., 1996), MoE training (Larochelle et al., 2009; Akbari et al., 2024), statistical convergence guarantees for classes of MoEs (Chen, 1995; Ho et al., 2022), robustness guarantees for such models (Puigcerver et al., 2022), amongst several other types of guarantees.
+However, to our knowledge, there are no available approximation guarantees for MoE or VC-dimension bounds for deep-learning-based MoEs.
+Thus, our results would be adding to the approximation theoretic foundations of MoE models as well as to the statistical foundations of deep-learning-based MoEs.
+Prototypes and Partitioning.
+Each region in our learned partition of the input space is associated with a representative point therein called a prototype.
+Prototypes (also called landmarks) are routinely used in image classification (Mensink et al., 2012), few-shot learning (Snell et al., 2017; Cao et al., 2020), dimensionality reduction (Law et al., 2019), in complex networks (Keller-Ressel & Nargang, 2023), and geometric deep learning (Ghadimi Atigh et al., 2021) to tractably encode massive structures.
+They are also standard in classical clustering algorithms such as $K$-medoids or $K$-means, wherein the part associated with each medoid (resp.
+centroid) defines a Voronoi cell or Voronoi region (Voronoi, 1908).
+Moreover, while partitioning is commonly employed in deep learning for various purposes, such as proving universal approximation theorems (Yarotsky, 2017; Lu et al., 2021b; Gühring & Raslan, 2021) or facilitating clustering-based learning (Zamanlooy & Kratsios, 2022; Trask et al., 2022; Ali & Nouy, 2023; Srivastava et al., 2022), existing approaches typically involve loading the entire model into VRAM.
+Our approach, however, differs by relying on a learned partition of the input space, where each part is associated with a distinct small neural network.
+Importantly, the complete set of networks forming the MoMLPs does not need to be simultaneously loaded into VRAM during training or inference.
+Paper Overview.
+Our paper is organized as follows.
+Section 3 contains preliminary notation, definitions, and mathematical background required for the formulation of our main results.
+Section 4 contains our main approximation (Theorem 4.1) and learning theoretic (Theorem 4.2) guarantees.
+Section 5 dives into the details of why MoEs can achieve arbitrary precision while maintaining a feasible active computational complexity by explaining the derivation of our main approximation theorem; the details of which are relegated to Appendix B.
+A technical version (Theorem 5.3) of our main approximation guarantee is then presented, which allows for the approximation of continuous functions of arbitrarily low regularity and for the organization of the experts defining the MoMLP into more general tree structures.
+Appendix D contains technical derivations of our main results as well as experimental elucidation of the benefit of MoEs, and specifically the toy MoMLP model.
+3 Preliminaries
+We standardize our notation, define the necessary mathematical formalisms to state our main results and define our toy MoE Model.
+3.1 Notation.
 We use the following notation: for any $f,g:\mathbb{R}\to\mathbb{R}$, we write $f\in\mathcal{O}(g)$ if there exist $x_{0}\in\mathbb{R}$ and $M\geq 0$ such that for each $x\geq x_{0}$ we have $|f(x)|\leq Mg(x_{0})$.
@@
 For each $n\in\mathbb{N}_{+}$ and $C\subseteq\mathbb{R}^{n}$, the indicator function $I_{C}$ of $C$ is defined by: for each $x\in\mathbb{R}^{n}$ set $I_{C}(x)=1$ if $x\in C$ and is 00 otherwise.
-Multi-Layer Perceptrons. In this paper, we will consider MLPs with trainable PReLU activation functions.
+3.2 Background
+Multi-Layer Perceptrons
+We will consider MLPs with trainable PReLU activation functions.
 Definition 3.1 (Trainable PReLU).
@@
 We identify any vector $\theta\in\mathbb{R}^{P([d])}$ with
-$\theta$ $\leftrightarrow\big{(}A^{(j)},b^{(j)},\bar{\alpha}^{(j)}\big{)}_{j=0}^{J}$ (3)
-$(A^{(j)},b^{(j)},\bar{\alpha}^{(j)})$ $\in\mathbb{R}^{d_{j+1}\times d_{j}}\times\mathbb{R}^{d_{j}}\times\mathbb{R}^{d_{j}}$
-With the above identification, we recursively define the representation function of a $[d]$-dimensional network by
+$\theta$ $\leftrightarrow\big{(}A^{(j)},b^{(j)},\bar{\alpha}^{(j)}\big{)}_{j=0}^{J}\,\,\mbox{ and }\,\,(A^{(j)},b^{(j)},\bar{\alpha}^{(j)})$ $\in\mathbb{R}^{d_{j+1}\times d_{j}}\times\mathbb{R}^{d_{j}}\times\mathbb{R}^{d_{j}}.$ (1)
+We recursively define the representation function of a $[d]$-dimensional network by
 $\mathbb{R}^{P([d])}\times\mathbb{R}^{d_{0}}\ni(\theta,x)$ $\mapsto\hat{f}_{\theta}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}A^{(J)}\,x^{(J)}+b^{(J)},$ (2)
-$x^{(j+1)}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sigma_{\bar{\alpha}^{(j)}}\bullet(A^{(j)}\,x^{(j)}+b^{(j)})$
+$x^{(j+1)}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sigma_{\bar{\alpha}^{(j)}}\bullet(A^{(j)}\,x^{(j)}+b^{(j)})\qquad\mbox{ for }j=0,\dots,J-1$
 $x^{(0)}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}x.$
-where $j$ runs from 00 to $J-1$.
 We denote by $\mathcal{NN}_{[d]}^{\sigma}$ the family of $[d]$-dimensional multilayer perceptrons (MLPs), $\{\hat{f}_{\theta}\}_{\theta\in\mathbb{R}^{P([d])}}$ described by (2).
@@
 The set of ReLU MLPs with depth $J$ and width $W$ is denoted by $\mathcal{NN}_{J,W:n,m}^{\sigma}=\cup_{[d]}\,\mathcal{NN}_{[d]}^{\sigma}$, where the union is taken over all multi-indices $[d]=[d_{0},\dots,d_{\tilde{J}}]$ with $n=d_{0}$, $m=d_{J+1}$, $d_{0},\dots,d_{J+1}\leq W$, and $\tilde{J}\leq J$.
-We now formally define the classes of neural pathways.
-Definition 3.2 (Neural Pathways).
-Let $J,W,L,n\in\mathbb{N}$ and fix an activation function $\sigma\in C(\mathbb{R})$.
-The set of neural pathways with at-most $L$ leaves, depth $J$, and width $W$, denoted by $\mathcal{NP}_{J,W,L:n,m}^{\sigma}$, consists of all functions $\hat{f}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ satisfying $\hat{f}=\sum_{i=1}^{l}\,f_{i}\,I_{C_{i}}$ where for $f_{1},\dots,f_{l}\in\mathcal{NN}^{\operatorname{PReLU}}_{J,W}$, and distinct prototypes $p_{1},\dots,p_{L}\in\mathbb{R}^{n}$; whose induced Voronoi cells $\{C_{i}\}_{i=1}^{L}$ are defined recursively by
-$C_{i}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{C}_{i}\setminus\bigcup_{j<i}\,\tilde{C_{j}}$ (5)
-where for $i=1,\dots,l$ the (non-disjoint) cells are
-$\tilde{C}_{i}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{\{}x\in[0,1]^{n}:\,\|x-p_{i}\|=\min_{j\in\{1,\dots,l\}}\,\|x-p_{j}\|\big{\}}$
-Remark 3.3 (Trees).
-The structure in any neural pathway is any tree with root node $\mathbb{R}^{n}$ and leaves given by the pairs $\{(p_{i},f_{i})\}_{i=1}^{l}$ or equivalently $\{(C_{i},f_{i})\}_{i=1}^{l}$.
-The purpose of any such tree is simply to efficiently route an input $x\in\mathbb{R}^{n}$ to one of the $l$ “leaf networks” $f_{1},\dots,f_{l}$ using $\mathcal{O}(l\log(l))$ queries; to identify which Voronoi cells $\{C_{i}\}_{i=1}^{l}$ the point $x$ is contained in.
-The partitioning technique used to define Eq.
-(5) is standard, see e.g.
-Krauthgamer et al.
-(2005, Proof of Lemma 1.7; page 846).
-It is employed to ensure disjointness of the Voronoi cells; this guarantees that no input is assigned to multiple prototypes.
-To keep notation tidy, we use $\mathcal{NP}_{J,W,L}^{\sigma}$ (resp.
-$\mathcal{N}^{\sigma}_{W,L}$), to abbreviate $\mathcal{NP}_{J,W,L:n,m}^{\sigma}$ (resp.
-$\mathcal{N}^{\sigma}_{W,L:n,m}$) whenever $n$ and $m$ are clear from the context.
 VC dimension
@@
 $\{I_{(0,\infty)}\circ f:\,f\in\mathcal{F}\}$ (3)
-shatters it in the classical sense; here $I_{(0,\infty)}(t)=1$ if $t\geq 0$ and equals to 00 otherwise.
+shatters it, i.e.
+if all possible classifiers on $\{x_{i}\}_{i=1}^{k}$ are implementable in the sense that $\{I_{(0,\infty)}\circ f:\,f\in\mathcal{F}\}=\{0,1\}^{\{x_{i}\}_{i=1}^{k}}$; here $I_{(0,\infty)}(t)=1$ if $t>0$ and equals to 00 otherwise.
 Denoted by $\operatorname{VC}(\mathcal{F})$, the VC dimension of $\mathcal{F}$ is the cardinality of the largest $k$-point subset shattered by $\mathcal{F}$.
@@
 Nearly matching upper bounds are in Bartlett et al. (2019).
-3.2 Main Theoretical Results
-To reduce notation, our main result is formulated for locally-Hölder function on the closed unit ball of $\mathbb{R}^{n}$, defined by $\overline{B}_{n}(0,1)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{x\in\mathbb{R}^{n}:\,\|x\|\leq 1\}$.
-Theorem 3.4 (Neural Pathways Avoid The Curse of Dimensionality).
-Fix $0<\alpha\leq 1$, dimensions $n,m\in\mathbb{N}$, parameters $0<\alpha\leq 1$ and an $\alpha$-Hölder map $f:\overline{\operatorname{B}}_{n}(0,1)\to\mathbb{R}^{m}$ with optimal Hölder coefficient $L\geq 0$, and $\sigma\in C(\mathbb{R})$ satisfying Definition 3.1.
-Fix an “approximation error” $\varepsilon>0$.
-There is a binary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq\overline{\operatorname{B}}_{n}(0,1)\times\mathbb{R}^{P([d])}$ satisfying
-$\max_{x\in\overline{\operatorname{B}}_{n}(0,1)}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|\in\mathcal{O}\big{(}\varepsilon^{1/\alpha-1/n}\big{)}.$ (8)
-Furthermore, for each $x\in\overline{\operatorname{B}}_{n}(0,1)$, if $v_{i}$ is such that $\|x-v_{i}\|\in\mathcal{O}(\varepsilon^{1/\alpha-1/n})$ then
-$\|f(x)-\hat{f}_{\theta_{i}}(x)\|<\varepsilon.$ (9)
-The depth and width of each $\hat{f}_{\theta_{i}}$ are recorded in Table 1 and the binary tree $\mathcal{T}$ has $2^{\mathcal{O}(n-\frac{n}{\alpha}\log(\varepsilon))}$  leaves.
-Remark 3.5.
-A technical version of Theorem 3.4 with $\nu$-ary trees and general uniformly continuous target functions, namely Theorem A.5, can be found in Appendix A.2.
-Next, we demonstrate that the neural pathways model can generalize and generate functions that are PAC-learnable, thanks to its finite VC dimension.
+We now define our toy mixture of experts model, the MoMLP.
+3.3 Definition: Our Toy Mixture of Experts Model - the MoMLP
+Figure 1: $1$) The distance from each input $x$ to all prototypes $p_{1},\dots,p_{8}$ ($\ell=8$) is queried.
+$2$) The network ($\hat{f}_{2}$ in the figure) assigned to the nearest prototype ($p_{2}$), is loaded onto the GPU and used for prediction.
+We study the following toy MoE model, where each expert (P)ReLU MLP specializes in a distinct region of the input domain $[0,1]^{n}$.
+Informally, these regions $C_{1},\dots,C_{\ell}$ correspond to the sets of closest points (Voronoi cells) from a finite set of prototypes/landmarks $p_{1},\dots,p_{\ell}$ in $[0,1]^{n}$, as illustrated in Figure 1.
+Associated to each region $C_{i}$ is a single expert MLP $\hat{f}_{i}$ with (P)ReLU activation function responsible for approximating the target function only thereon.
+Here, the sparse gating procedure which routes any given input $x\in[0,1]^{n}$ to the expert corresponding to the nearest prototype $p_{i}$ is implemented by a (finite) routing tree $\mathcal{T}=(V,E)$ whose nodes $V$ are points in $[0,1]^{n}$ and leaves (terminal nodes) are the points $p_{1},\dots,p_{\ell}$.
+We now formally define the classes of MoMLPs.
+Definition 3.2 (MoMLPs).
+Let $J,W,{l},n\in\mathbb{N}$ and fix an activation function $\sigma\in C(\mathbb{R})$.
+The set of MoMLPs with at-most $L$ leaves, depth $J$, and width $W$, denoted by $\mathcal{NP}_{J,W,{l}:n,m}^{\sigma}$, consists of all functions $\hat{f}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ satisfying $\hat{f}=\sum_{i=1}^{l}\,f_{i}\,I_{C_{i}}$ where for $f_{1},\dots,f_{l}\in\mathcal{NN}^{\operatorname{PReLU}}_{J,W}$, and distinct prototypes $p_{1},\dots,p_{{l}}\in\mathbb{R}^{n}$; whose induced Voronoi cells $\{C_{i}\}_{i=1}^{{l}}$ are defined recursively by
+$C_{i}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{C}_{i}\setminus\bigcup_{j<i}\,\tilde{C_{j}},$ (5)
+where for $i=1,\dots,l$ the (non-disjoint) cells are
+$\tilde{C}_{i}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{\{}x\in[0,1]^{n}:\,\|x-p_{i}\|=\min_{j\in\{1,\dots,l\}}\,\|x-p_{j}\|\big{\}}.$ (6)
+The Routing Trees
+The structure in any MoMLP is any tree with root node $\mathbb{R}^{n}$ and leaves given by the pairs $\{(p_{i},f_{i})\}_{i=1}^{l}$ or equivalently $\{(C_{i},f_{i})\}_{i=1}^{l}$.
+The purpose of any such tree is simply to efficiently route an input $x\in\mathbb{R}^{n}$ to one of the $l$ “leaf networks” (the experts) $f_{1},\dots,f_{l}$ using $\mathcal{O}(l\log(l))$ queries; to identify which Voronoi cells $\{C_{i}\}_{i=1}^{l}$ the point $x$ is contained in.
+Remark 3.3 (Partitioning in (5) in classical computer science).
+The partitioning technique used to define Eq.
+(5) is standard, see e.g.
+Krauthgamer et al.
+(2005, Proof of Lemma 1.7; page 846).
+It is employed to ensure disjointness of the Voronoi cells; this guarantees that no input is assigned to multiple prototypes.
+To keep notation tidy, we use $\mathcal{NP}_{J,W,L}^{\sigma}$ (resp.
+$\mathcal{N}^{\sigma}_{W,L}$), to abbreviate $\mathcal{NP}_{J,W,L:n,m}^{\sigma}$ (resp.
+$\mathcal{N}^{\sigma}_{W,L:n,m}$) whenever $n$ and $m$ are clear from the context.
+4 Main Results
+We first present our main approximation theoretic guarantee, which gives complexity estimates for mixtures of MLPs with trainable PReLU activation functions when uniformly approximating arbitrary locally-Hölder function on the closed unit ball of $\mathbb{R}^{n}$, defined by $\overline{B}_{n}(0,1)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{x\in\mathbb{R}^{n}:\,\|x\|\leq 1\}$.
+Our rates depend on a “number of experts-to-expert complexity trade-off parameter” $r\in\mathbb{R}$ which determines how fast the overall MoE complexity scales, in terms of the number of experts and the complexity of each expert, as the approximation error becomes small.
+Setting $r<0$ implies that more model complexity will be loaded into each expert MLP and there will be fewer experts defining the MoE.
+In contrast, setting $r>0$ loads less complexity in each expert MLP at the cost of more experts in the MoE.
+In particular, when $r=0$, each expert will have constant complexity even when the approximation error becomes arbitrarily small.
+Theorem 4.1 (Trade-Off: No. Expert vs. Expert Complexity).
+Suppose that $\sigma$ satisfies Definition 3.1. Fix an “number of experts-to-expert complexity trade-off parameter” $r\in\mathbb{R}$.
+For every $\alpha$-Hölder map $f:\overline{B}_{n}(0,1)\to\mathbb{R}^{m}$ with $0<\alpha\leq 1$ and each approximation error $\varepsilon>0$, there is a $p\in\mathbb{N}_{+}$, a binary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq{\mathcal{K}}\times\mathbb{R}^{p}$ and a family of MLPs with (P)ReLU activation function $\{\hat{f}_{\theta_{i}}\}_{i=1}^{L}$ defined by $p$ parameters and mapping $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$ satisfying:
+$\max_{x\in{\mathcal{K}}}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|\in\Theta\big{(}\varepsilon^{1/\alpha-2r/n}\big{)}$
+and for each $x\in{\mathcal{K}}$ and $i=1,\dots,L$, if $\|x-v_{i}\|<\delta$ then
+$\|f(x)-f_{\theta_{i}}(x)\|<\varepsilon.$
+The depth and width of each $\hat{f}_{\theta_{i}}$ and the number of leaves, height, and number of nodes required to build the $\nu$-ary tree are all recorded in Table 1.
+Next, we demonstrate that the MoMLP model can generalize and generate functions that are PAC-learnable, thanks to its finite VC dimension.
 This property, however, breaks down in MLP models with super-expressive activation functions.
-Proposition 3.6 (Neural Pathways Can Generalize).
+Theorem 4.2 (VC-Dimension Bounds for MoMLPs - MoMLPs Can Generalize).
 Let $J,W,L,n\in\mathbb{N}_{+}$. Then $\operatorname{VC}\big{(}\mathcal{NN}_{J,W,L:n,1}^{\operatorname{PReLU}}\big{)}$ is of
-$\mathcal{O}\big{(}L\log(L)^{2}\,\max\{nL\log(L),JW^{2}\log(JW)\}\big{)}$
+$\mathcal{O}\big{(}L\log(L)^{2}\,\max\{nL\log(L),JW^{2}\log(JW)\}\big{)}$ (7)
 In particular, $\mathrm{VC}(\mathcal{NP}_{J,W,L:n,1}^{ReLU})<\infty$.
-We complement the main result of Bartlett et al.
-(2019) by demonstrating that the class of unstable MLPs (Shen et al., 2022a) possesses infinite VC dimension, even when they have finite depth and width.
-Thus, while they may serve as a gold standard from the perspective of approximation theory, they should not be considered a benchmark gold standard from the viewpoint of learning theory.
-Proposition 3.7 (MLPs with Super-Expressive Activation Do Not Generalize).
-Let $\mathcal{F}$ denote the set of MLPs with activation function in Definition A.1, depth at-most $15$, and width at-most $36n(2n+2)$. Then $\operatorname{VC}(\mathcal{F})=\infty$.
-The VC dimension bounds for the standard MLP model, MLP with a super-expressive activation function as proposed by Shen et al. (2022a), and the neural pathways model are summarized in Table 2.
-Table 2: VC Dimension of the Neural Pathways, ReLU MLP, and MLP model with Super-Expressive Activation function of Shen et al.
-(2022a).
-All models have depth $J$, width $W$, and (when applicable) $L$ leaves; where $J,W,L,n\in\mathbb{N}_{+}$.
-Model VC Dim Ref.
-Neural Pathways $\mathcal{O}\big{(}L\log(L)^{2}\,\max\{nL\log(L),JW^{2}\log(JW)\}\big{)}$ Prop 3.6 ReLU MLP $\mathcal{O}\big{(}JW^{2}\log(JW)\big{)}$ Bartlett et al.
-(2019) Super-Expressive $\infty$ Prop 3.7
-4 Related Work and Discussion
-With the advent of large foundation models, the use of distributed computing to scale deep learning models beyond what is feasible to process on a single machine has become exceedingly relevant.
-However, most of the literature focuses on distributed or parallelized GPU training of standard neural network models (Dean et al., 2012; Hegde & Usmani, 2016; Sergeev & Del Balso, 2018; Ben-Nun & Hoefler, 2019).
-Next, we review relevant literature.
-Prototypes.
-Each region in our learned partition of the input space is associated with a representative point therein called a prototype.
-Prototypes (also called landmarks) are routinely used in image classification (Mensink et al., 2012), few-shot learning (Snell et al., 2017; Cao et al., 2020), dimensionality reduction (Law et al., 2019), in complex networks (Keller-Ressel & Nargang, 2023), and geometric deep learning (Ghadimi Atigh et al., 2021) to tractably encode massive structures.
-They are also standard in classical clustering algorithms such as $K$-medoids or $K$-means, wherein the part associated with each medoid (resp.
-centroid) defines a Voronoi cell or Voronoi region (Voronoi, 1908).
-In our proposed approach, one key difference is that we optimize prototypes to maximize performance on the loss function defining the learning problem, rather than to divide up the input space based on a clustering objective, which may not optimize the model’s performance.
-Nevertheless, in principle, many of these prototype-based clustering algorithms can be substituted into our pipeline.
-Partitioning.
-While partitioning is commonly employed in deep learning for various purposes, such as proving universal approximation theorems (Yarotsky, 2017; Lu et al., 2021b; Gühring & Raslan, 2021) or facilitating clustering-based learning (Kratsios & Zamanlooy, 2022; Trask et al., 2022; Ali & Nouy, 2023; Srivastava et al., 2022), existing approaches typically involve loading the entire model into VRAM.
-Our approach, however, differs by relying on a learned partition of the input space, where each part is associated with a distinct small neural network.
-Importantly, the complete set of networks forming the neural pathways does not need to be simultaneously loaded into VRAM during training or inference.
-Cluster-then-Predict.
-Several deep learning models employ a cluster-then-predict approach to upscaling, e.g., (Wang et al., 2021).
-In these approaches, one typically maps an input to a specific prototype and passes that prototype through a relatively small network, effectively reducing dimension by discretizing the input space to consist only of the set of prototypes.
-This is in stark contrast to the neural pathways framework, which performs no discretization and only hierarchically partitions the input space.
-Every input is passed through a neural network at the leaf node, corresponding to the nearest prototypes, but the input itself is never replaced by one of the prototypes.
-Srivastava et al.
-(2022) also use multiple local networks, but they employ soft clustering for the final prediction, which means they have to load all the local models for a single sample.
-Comparison to Standard Distributed Training.
-One can distribute the complexity of feedforward models by storing each of their layers in offline memory and then loading them sequentially into VRAM during the forward pass.
-This does avoid loading more than $\mathcal{O}(\operatorname{Width})$ active parameters into VRAM at any given time, where $\operatorname{Width}$ denotes the width of the feedforward model.
-However, doing so implies that all the model parameters are ultimately loaded during the forward pass.
-This contrasts with the neural pathways model, which requires $\mathcal{O}\big{(}\log_{2}(K)\,\operatorname{Width}^{2}\,\operatorname{Depth}\big{)}$ to be loaded into memory during a forward pass; where $\operatorname{Width}$ and $\operatorname{Depth}$ are respectively the largest width and depth of the MLP at any leaf of the tree defining a given neural pathway, and $K$ denotes the number of prototypes.
-However, in the forward pass, one loads $\mathcal{O}(\varepsilon^{-n/2})$ parameters for the best worst-case MLP while only $\mathcal{O}(n\log(1/\varepsilon)/\varepsilon)$ are needed in the case of the neural pathways.
-The number of parameters here represents the optimal worst-case rates for both models (see Table 1 and Theorem A.5).
+4.1 Discussion
+Trade-off between Number of Experts and Expert Complexity.
+Our results suggest that, theoretically, successful MoE models may not need each expert to be highly overparameterized if there are enough experts.
+This hypothesis is ablated experimentally in Appendix D in the context of irregular function approximation in low-dimension space; which is equivalent to high-dimensional regular function approximation (see Appendix C for a discussion on this later point).
 Pruning.
@@
 Logarithmic number of queries via trees.
-For many prototypes, as in our main guarantee, the neural pathways only need to evaluate the distance between the given input and a logarithmic number of prototypes—specifically, one for each level in the tree—when using deep binary trees to hierarchically refine the Voronoi cells.
-Thus, a given machine never processes the exponential number of prototypes, and only $\nu\lceil\log_{\nu}(K)\rceil$ prototypes are ever queried for any given input; when trees are $\nu$-ary (as in Theorem A.5), and where $K$ denotes the number of prototypes.
-Since we consider that prototypes are queried separately and before loading neural pathways, we do not take them into account when counting the number of learnable parameters.
+For many prototypes, as in our main guarantee, the MoMLPs only need to evaluate the distance between the given input and a logarithmic number of prototypes—specifically, one for each level in the tree—when using deep binary trees to hierarchically refine the Voronoi cells.
+Thus, a given machine never processes the exponential number of prototypes, and only $\nu\lceil\log_{\nu}(K)\rceil$ prototypes are ever queried for any given input; when trees are $\nu$-ary (as in Theorem 5.3), and where $K$ denotes the number of prototypes.
+Since we consider that prototypes are queried separately and before loading MoMLPs, we do not take them into account when counting the number of learnable parameters.
 Moreover, the size of our prototypes is negligible in our experiments.
-5 Experiments
-We evaluate our approach in two standard machine learning tasks: regression and classification.
-We experimentally show that neural pathways, which distribute predictions over multiple neural networks, are competitive with a single large neural network containing as many model parameters as all the neural pathways combined.
-This is desirable in cases where the large neural network does not fit into the memory of a single machine.
-On the contrary, the neural pathways framework can be trained by distributing each neural pathway on separate machines (or equivalently serially on a single machine).
-Inference can then be performed by loading only a single neural pathway at a time into the GPU.
-Our source code to reproduce our results is available in the supplementary material.
-5.1 Regression
-We first consider regression, where the goal is to approximate non-convex synthetic functions often used for performance test problems.
-In particular, we choose 1-dimensional Hölder functions, as well as Ackley (Ackley, 1987) and Rastrigin (Rastrigin, 1974) functions, whose formulations are detailed in Appendix D.
-1D Hölder Functions.
-We illustrate our primary finding, as encapsulated in Theorem 3.4, by leveraging $1$-dimensional functions characterized by very low regularity.
-This choice is motivated by the jagged structure inherent in such functions, necessitating an exponentially higher sampling frequency compared to smooth functions for achieving an accurate reconstruction.
-Indeed, this crucial sampling step forms the foundation of many quantitative universal approximation theorems (Yarotsky, 2017; Shen et al., 2021; Kratsios & Papon, 2022).
-As elaborated in Appendix C, approximating a well-behaved (Lipschitz) function in $d$ dimensions poses a challenge equivalent to approximating a highly irregular function ($1/d$-Hölder) in a single dimension.
-A visual representation of a $1/d$-Hölder function is presented in Figure 2, exemplified by the trajectory of a fractional Brownian motion with a Hurst parameter of $\alpha=1/d$.
-A formal definition is available in Appendix C.
-Figure 2: Visual Comparison of Functions with High ($\alpha\approx 1$) vs.
-Low ($\alpha\approx 0$) Hölder regularity.
-If $\alpha\approx 1$, the function (green) is approximately differentiable almost everywhere, meaning it does not osculate much locally and thus is simple to approximate.
-If $\alpha\approx 0$, the function may be nowhere differentiable and jagged; its extreme details make it difficult to approximate.
-Table 3: Test mean squared error (average and standard deviation) for the different functions of the regression task.
-1D Hölder 2D Ackley 3D Ackley 2D Rastrigin 3D Rastrigin
-Neural pathways (ours) 0.057 $\pm$ 0.085 0.00015 $\pm$ 0.00006 0.00068 $\pm$ 0.00010 0.0480 $\pm$ 0.0073 1.0062 $\pm$ 0.0446
-Baseline 0.128 $\pm$ 0.012 0.08723 $\pm$ 0.01059 0.09303 $\pm$ 0.03156 3.0511 $\pm$ 0.3581 8.0376 $\pm$ 4.0499
-2D and 3D Functions. We select the Ackley and Rastrigin functions, with their respective 2D representations showcased in Figure 3, as widely recognized benchmarks in the field of optimization.
-(a) Ground Truth Ackley
-(b) Ground Truth Rastrigin
-(c) Predicted Ackley
-(d) Predicted Rastrigin
-(e) log MSE
-(f) log MSE
-Figure 3: Comparison of ground truth and predicted results for 2D Ackley and Rastrigin functions over the domain $[-1,1]^{2}$.
-Evaluation protocol.
-We consider the setup where the domain of a function that we try to approximate is the $n$-dimensional closed set $[a,b]^{n}$.
-For instance, we arbitrarily choose the domain $[0,1]$ when $n=1$, and $[-1,1]^{n}$ when $n\geq 2$.
-Our training and test samples are the $s^{n}$ vertices of the regular grid defined on $[a,b]^{n}$.
-At each run, $80\%$ of the samples are randomly selected for training and validation, and the remaining $20\%$ for testing.
-During training, for a given and fixed set of $K$ prototypes $p\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(p_{1},\dots,p_{K})$, we assign each training sample $x$ to its nearest prototype $p_{k}$ and associated neural network $\hat{f}_{k}$.
-We learn the prototypes as explained in Section 2.2.
-For simplicity, we set the number of prototypes to $K=4$; we also set $s=10,000$ if $n=1$, $s=150$ if $n=2$ (i.e., $500150^{2}=22,$ 500 samples), and $s=30$ if $n=3$ (i.e., $27,000$ samples).
-More details can be found in Appendix D.
-Test performance.
-In Table 3, we present the mean squared error obtained on the test set across 10 random initializations and various splits of the training/test sets.
-The baseline consists of a single neural network with the same overall architecture as each neural pathway but possesses as many parameters as all the neural pathways combined.
-In all instances, the neural pathways model demonstrates a significant performance advantage compared to the baseline.
-Figure 3 illustrates the predictions generated by our neural pathways, showcasing the capability of our approach to achieve a good approximation of the ground truth functions.
-5.2 Classification
-Datasets.
-We evaluate classification on standard image datasets such as CIFAR-10 (Krizhevsky & Hinton, 2010), CIFAR-100, and Food-101 (Bossard et al., 2014), which consist of 10, 100, and 101 different classes, respectively.
-We use the standard splits of training/test sets: the datasets include (per category) 5,000 training and 1,000 test images for CIFAR-10, 500 training and 100 test images for CIFAR-100, and 750 and 250 for Food-101.
-Training.
-Our neural pathways model takes as input latent DINOv2 encodings (Oquab et al., 2023) of images from the aforementioned datasets.
-Each sample $x\in\mathbb{R}^{768}$ corresponds to a DINOv2 embedding (i.e., $n=768$).
-Additionally, we set the prototypes as centroids obtained through the standard $K$-means clustering on the DINOv2 embedding space.
-The replacement of our original prototype learning algorithm is sensible in this context, as we operate within a structured latent space optimized through self-supervised learning using large-scale compute and datasets.
-Table 4: Test classification accuracy using DINOv2 features as input (average and standard deviation).
-Dataset CIFAR-10 CIFAR-100 Food-101
-Ours (weighted) 98.40 $\pm$ 0.05 90.01 $\pm$ 0.11 91.86 $\pm$ 0.10
-Ours (unweighted) 98.42 $\pm$ 0.04 89.62 $\pm$ 0.25 91.79 $\pm$ 0.16
-Baseline 98.45 $\pm$ 0.06 89.85 $\pm$ 0.17 91.45 $\pm$ 1.09
-Due to the potential class imbalance in the various Voronoi cells formed by the prototypes, we utilize two variations of the cross-entropy loss for each neural pathway $\hat{f}_{k}$.
-The first variation, termed unweighted, assigns equal weight to all categories.
-The weighted variation assigns a weight that is inversely proportional to the distribution of each category in the Voronoi cell defined by the prototype $p_{k}$.
-Test performance.
-We present the test classification accuracy over 10 random initializations of both the baseline and our neural pathways in Table 4.
-The weighted version performs slightly better on datasets with a large number of categories.
-Our approach achieves comparable results with the baseline, effectively decomposing the prediction across multiple smaller models.
-6 Conclusion
-We introduced a novel approach to overcome the curse of dimensionality based on distributed neural computation.
-Instead of relying on exceptionally regular target functions or super-expressive MLPs, as seen in other related works in the literature, we propose distributing computation across several small specialized networks.
-These networks can learn to model the target function within Voronoi cells of the input space.
-Our approach is particularly relevant, given the computational limitations of current AI hardware, handicapped by finite VRAM.
-This problem has recently been exacerbated by the growing trend to increase model size with large foundation models.
-In particular, our neural pathways framework can scale up simply by adding more neural networks (or equivalently by refining the Voronoi cells which subdivide the input space), since each independent network can always be loaded into VRAM while the overall model keeps increasing in capacity.
-We prove, via Theorem 3.4 and A.5, that the number of parameters of any single neural pathway loaded in active GPU memory only requires $\mathcal{O}(n/\varepsilon)$ depth and $\mathcal{O}(n)$ width when approximating a Lipschitz map over the cube $[0,1]^{n}$ uniformly to an accuracy of $\varepsilon>0$.
-This contrasts with the optimal approximation rates for standard (non-distributed) MLP architectures, which require $\mathcal{O}(\varepsilon^{-n})$ depth and $\mathcal{O}(n)$ width to achieve the same $\varepsilon$ accuracy.
-Importantly, unlike results that break the curse of dimensionality using standard MLPs, our results always hold in practice since we do not require the target function to have high regularity.
-Furthermore, in contrast to approximation results leveraging super-expressive activation functions with infinite VC-dimension (Proposition 3.7), any neural pathway of finite depth and width, defined by a finite tree with weight and bias constraints, has finite VC-dimension (Proposition 3.6).
-Thus, our models can learn and generalize effectively when trained on finite data.
-We validate our algorithmic framework and mathematical analysis by testing neural pathways on regression tasks.
-Specifically, we focus on learning 1D Hölder functions of low regularity, as well as on 2D and 3D Ackley and Rastrigin functions.
-Additionally, we assess its performance in classification, particularly in higher-dimensional spaces such as the latent space of pretrained self-supervised large computer vision models like DINOv2.
-This not only helps verify the applicability of our model beyond synthetic experiments but also demonstrates its effectiveness in setups that are of closer interest to the applied machine learning community.
-In all cases, our model either matches or outperforms the centralized baseline while loading substantially lighter networks into GPU memory one at a time.
-Future Work.
-In this paper, we present the initial implementation of our framework and we anticipate that future efforts will be necessary to scale it to larger networks.
-In such cases, the inclusion of deeper tree structures, akin to those discussed in our proofs, will be pertinent in addressing computational complexity.
-Similarly, as discussed in Section 2, our framework is modular, and we foresee that the methodology could be potentially modified depending on the downstream task and specific network architectures, as we primarily focused on MLPs only.
-7 Acknowledgment and Funding
-The authors would like to thank James Lucas and Rafid Mahmood for their helpful feedback while preparing the manuscript.
-AK acknowledges financial support from the NSERC Discovery Grant No. RGPIN-2023-04482. HSOB acknowledges financial support from an EV travel grant.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Overview of Technical Derivation
-We now overview the proof of our main result and its full technical formulation. These objectives require us to recall definitions from the analysis of metric spaces, which we do now.
-A.1 Preliminary Definitions
+4.2 Application: Controlling The Complexity in VRAM maintaining a Finite VC Dimension
+Super-Expressive Activation Functions Have Infinite VC-Dimension.
+We complement the main result of Bartlett et al.
+(2019) by demonstrating that the class of unstable MLPs (Shen et al., 2022a) possesses infinite VC dimension, even when they have finite depth and width.
+Thus, while they may serve as a gold standard from the perspective of approximation theory, they should not be considered a benchmark gold standard from the viewpoint of learning theory.
+We consider a mild extension of the super-expressive activation function of Shen et al.
+(2022a).
+This parametric extension allows it to implement the identity map on the real line as well as the original super-expressive activation function thereof.
+Definition 4.3 (Trainable Super-Expressive Activation Function).
+A trainable action function $\sigma:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is of super-expressive type if for all $\alpha\in\mathbb{R}$
+$\sigma_{\alpha}:\mathbb{R}\ni x\mapsto\alpha x+(1-\alpha)\sigma^{\star}(x)$
+where $\sigma^{\star}:\mathbb{R}\to\mathbb{R}$ is given by
+$\sigma^{\star}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\begin{cases}|x\mod{2}|&\mbox{ for }0\leq x\\ \frac{x}{|x|+1}&\mbox{ for }x<0.\end{cases}$ (8)
+Proposition 4.4 (MLPs with Super-Expressive Activation Do Not Generalize).
+Let $\mathcal{F}$ denote the set of MLPs with activation function in Definition 4.3, depth at-most $15$, and width at-most $36n(2n+2)$. Then $\operatorname{VC}(\mathcal{F})=\infty$.
+The VC dimension bounds for the standard MLP model, MLP with a super-expressive activation function as proposed by Shen et al. (2022a), and the MoMLP model are summarized in Table 2.
+Table 2: VC Dimension of the MoMLPs, ReLU MLP, and MLP model with Super-Expressive Activation function of Shen et al.
+(2022a).
+All models have depth $J$, width $W$, and (when applicable) $L$ leaves; where $J,W,L,n\in\mathbb{N}_{+}$.
+Model VC Dim Ref.
+MoMLPs $\mathcal{O}\big{(}L\log(L)^{2}\,\max\{nL\log(L),JW^{2}\log(JW)\}\big{)}$ Thrm 4.2
+ReLU MLP $\mathcal{O}\big{(}JW^{2}\log(JW)\big{)}$ Bartlett et al. (2019)
+Super-Expressive $\infty$ Prop 4.4
+5 Overview of Derivation
+We now overview the proof of our main result and its full technical formulation.
+These objectives require us to recall definitions from the analysis of metric spaces, which were not required in the statement of our main result but which are required when overviewing our proof.
+5.1 Technical Definitions
 The metric ball in $(\mathcal{X},d)$ of radius $r>0$ at $x\in\mathcal{X}$ is denoted by $\operatorname{Ball}_{(\mathcal{X},d)}(x,r)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{z\in\mathcal{X}\colon d(x,z)<r\}$.
-The metric space $(\mathcal{X},d)$ is called doubling, if there is $C\in\mathbb{N}_{+}$ for which every metric ball in $(\mathcal{X},d)$ can be covered by at most $C$ metric balls of half its radius.
+A metric space $(\mathcal{X},d)$ is called doubling, if there is $C\in\mathbb{N}_{+}$ for which every metric ball in $(\mathcal{X},d)$ can be covered by at most $C$ metric balls of half its radius.
 The smallest such constant is called $(\mathcal{X},d)$’s doubling number, and is here denoted by $C_{(\mathcal{X},d)}$.
-Though this definition may seem abstract at first, (Heinonen, 2001, Theorem 12.1), provides an almost familiar characterization of all doubling metric spaces; indeed, $\mathcal{K}$ is a doubling metric space if and only if it can be identified via a suitable invertible map [^1] So called quasi-symmetric maps, see (Heinonen, 2001, page 78).
+Though this definition may seem abstract at first, Heinonen (2001, Theorem 12.1) provides an almost familiar characterization of all doubling metric spaces; indeed, $\mathcal{K}$ is a doubling metric space if and only if it can be identified via a suitable invertible map [^5] So called quasi-symmetric maps, see (Heinonen, 2001, page 78).
 with a subset of some Euclidean space.
-Every subset of $\mathbb{R}^{n}$, for any $n\in\mathbb{N}_{+}$, is a doubling metric space; see (Robinson, 2011, Chapter 9) for details.
+Every subset of $\mathbb{R}^{n}$, for any $n\in\mathbb{N}_{+}$, is a doubling metric space; see Robinson (2011, Chapter 9) for details.
 Example 1 (Subsets of Euclidean Spaces).
 Fix a dimension $n\in\mathbb{N}_{+}$.
-The doubling number of any subset of Euclidean space is [^2] See (Robinson, 2011, Lemma 9.2) and the brief computations made in the proof of (Robinson, 2011, Lemma 9.4).
+The doubling number of any subset of Euclidean space is [^6] See Robinson (2011, Lemma 9.2) and the brief computations made in the proof of Robinson (2011, Lemma 9.4).
 at most $2^{n+1}$.
@@
 $\log_{v}$ is base $v$ for a given $v\in\mathbb{N}_{+}$ and $\log\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log_{2}$.
-As on (Petrova & Wojtaszczyk, 2023, page 762), the radius of a subset $A\subseteq\mathbb{R}^{n}$, denoted by $\operatorname{rad}(A)$, is defined by
-$\operatorname{rad}(A)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\inf_{x\in\mathbb{R}^{n}}\,\sup_{a\in A}\,\|x-a\|.$
+As in Petrova & Wojtaszczyk (2023, page 762), the radius of a subset $A\subseteq\mathbb{R}^{n}$, denoted by $\operatorname{rad}(A)$, is defined by
+$\operatorname{rad}(A)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\inf_{x\in\mathbb{R}^{n}}\,\sup_{a\in A}\,\|x-a\|.$ (9)
 The diameter of any such set $A$, denoted by $\operatorname{diam}(A)$, satisfies the sharp inequality $\operatorname{diam}(A)\leq 2\operatorname{rad}(A)$.
@@
 Furthermore, we note that every $(\alpha,L)$-Hölder function is uniformly continuous with modulus of continuity $\omega(t)=L\,t^{\alpha}$.
-We consider a mild extension of the super-expressive activation function of Shen et al.
-(2022a).
-This parametric extension allows it to implement the identity map on the real line as well as the original super-expressive activation function thereof.
-Definition A.1 (Trainable Super-Expressive Activation Function).
-A trainable action function $\sigma:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is of super-expressive type if for all $\alpha\in\mathbb{R}$
-$\sigma_{\alpha}:\mathbb{R}\ni x\mapsto\alpha x+(1-\alpha)\sigma^{\star}(x)$
-where $\sigma^{\star}:\mathbb{R}\to\mathbb{R}$ is given by
-$\sigma^{\star}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\begin{cases}|x\mod{2}|&\mbox{ for }0\leq x\\ \frac{x}{|x|+1}&\mbox{ for }x<0.\end{cases}$ (10)
-A.2 Proof Sketch
-Lemma A.2.
+5.2 Helping to Explain MoEs via Proof Sketch
+Lemma 5.1 (Size of a Tree Whose Nodes Form a $\delta$-net of a Compact Subset of $\mathbb{R}^{n}$).
 Let $\mathcal{K}$ be a compact subset of $\mathbb{R}^{n}$ whose doubling number is $C$.
@@
 where $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}1/\log(v)$.
-At each node of the tree, we will place an MLP which only locally approximates the target function on a little ball of suitable radius (implied by the tree valency $v$ and height $h$) of lemma A.2.
+At each node of the tree, we will place an MLP which only locally approximates the target function on a little ball of suitable radius (implied by the tree valency $v$ and height $h$) of lemma 5.1.
 I.e.
-by the storage space we would like to allocate to our neural pathway model.
+by the storage space we would like to allocate to our MoMLP model.
 The next step of the proof relies on a mild extension of the quantitative universal approximation theorem in Shen et al.
@@
 (2023, Proposition 3.10) beyond the Hölder case.
-Lemma A.3 (Universal Approximation Theorem).
-Let $n,m\in\mathbb{N}_{+}$ with $n\geq 3$, ${\mathcal{K}}\subseteq\mathbb{R}^{n}$ be compact set $f:{\mathcal{K}}\to\mathbb{R}^{m}$ be uniformly continuous with strictly monotone continuous modulus of continuity $\omega$.
-Let $\sigma$ be an activation function as in Definitions A.1 or 3.1.
+Lemma 5.2 (Vector-Valued Universal Approximation Theorem with Explicit Diameter Dependence).
+Let $n,m\in\mathbb{N}_{+}$ with $n\geq 3$, ${\mathcal{K}}\subseteq\mathbb{R}^{n}$ be compact set of radius $\delta\geq 0$, $f:{\mathcal{K}}\to\mathbb{R}^{m}$ be uniformly continuous with strictly monotone continuous modulus of continuity $\omega$.
+Let $\sigma$ be an activation function as in Definitions 4.3 or 3.1.
 For each $\varepsilon>0$, there exists an MLP $\hat{f}_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with trainable activation function $\sigma$ satisfying the uniform estimate
@@
 The depth and width of $\hat{f}$ are recorded in Table 3.
-Table 5: Complexity of the MLP $\hat{f}_{\theta}$ in Lemma A.3.
-Activation $\sigma$ Super Expressive A.1 PReLU 3.1
-Depth $(J)$ $15m$ $m\left(19+2n+11\Big{\lceil}\Big{(}\frac{\operatorname{rad}(K)2^{3/2}n^{1/2}}{(n+1)^{1/2}\omega^{-1}\big{(}\varepsilon/(131\sqrt{n\,m})\big{)}}\Big{)}^{n/2}\Big{\rceil}\right)$
-Width ($\max_{j}\,d_{j}$) $36n(2n+1)+m$ $16\max\{n,3\}+m$
-Combining Lemmata A.2 and A.3 we obtain Theorem 3.4.
-Lemma A.4.
-Let ${\mathcal{K}}$ be a compact subset of $\mathbb{R}^{n}$ whose doubling number is $C$ and a uniformly continuous map $f:{\mathcal{K}}\to\mathbb{R}^{m}$ with modulus of continuity $\omega$.
-Fix $v\in\mathbb{N}$ with $v\geq 2$, $0<\delta\leq\operatorname{rad}({\mathcal{K}})$, and $\varepsilon>0$.
-Suppose that $\sigma$ satisfies Definition 3.1.
-There exists a $p\in\mathbb{N}_{+}$ and a $v$-ary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq{\mathcal{K}}\times\mathbb{R}^{p}$ satisfying
-$\max_{x\in{\mathcal{K}}}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|<\delta.$ (12)
-Furthermore, for each $x\in{\mathcal{K}}$ and each $i=1,\dots,L$, if $\|x-v_{i}\|<\delta$ then
-$\|f(x)-f_{\theta_{i}}(x)\|<\varepsilon.$
-We have the following estimates:
-  1. (i)
-    Depth: Depth of each $\hat{f}_{\theta_{i}}$ is $m\left(19+2n+11\Big{\lceil}\Big{(}\frac{\delta 2^{3/2}n^{1/2}}{(n+1)^{1/2}\omega^{-1}\big{(}\varepsilon/(131\sqrt{n\,m})\big{)}}\Big{)}^{n/2}\Big{\rceil}\right)$
-  2. (ii)
-    Width: Width of each $\hat{f}_{\theta_{i}}$ is $16\,\max\{n,3\}+m$
-  3. (iii)
-    Leaves: at most $L=v^{\big{\lceil}c\,\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil},}$
-  4. (iv)
-    Height: $\big{\lceil}c\,\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil}$,
-  5. (v)
-    Nodes: At most $\frac{v^{\lceil c\,\log(C)(1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}})))\rceil+1}-1}{\big{\lceil}c\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil}-1}$
-where $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
+Table 3: Complexity of the MLP $\hat{f}_{\theta}$ in Lemma 5.2. See Table 5 in Appendix A for more detailed estimates.
+Activation $\sigma$ Super Expressive 4.3 PReLU 3.1
+Depth $(J)$ $\mathcal{O}(1)$ $\mathcal{O}\big{(}(\delta/\omega^{-1}(\varepsilon))^{n/2}\big{)}$
+Width ($\max_{j}\,d_{j}$) $\mathcal{O}(1)$ $\mathcal{O}(1)$
+Combining Lemmata 5.1 and 5.2 we obtain Theorem 4.1.
 We now present the technical version of Theorem 4.1.
 This result allows distributed neural computing using $\nu$-ary trees and allows for the approximation general uniformly continuous target functions.
-Theorem A.5 (Technical Version of Theorem 3.4).
-Let ${\mathcal{K}}$ be a compact subset of $\mathbb{R}^{n}$ whose doubling number is $C$ and a uniformly continuous map $f:{\mathcal{K}}\to\mathbb{R}^{m}$ with modulus of continuity $\omega$.
-Fix an “approximation rate” $r\geq 0$, a “valency parameter” $v\in\mathbb{N}$ with $v\geq 2$, and an “approximation error” $\varepsilon>0$.
+Theorem 5.3 (Trade-Off: No. Expert vs. Expert Complexity - Technical Version of Theorem 4.1).
 Suppose that $\sigma$ satisfies Definition 3.1.
-Then, there is an $v$-ary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq{\mathcal{K}}\times\mathbb{R}^{p}$ satisfying
-$\max_{x\in{\mathcal{K}}}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|<\frac{\varepsilon^{-2r/n}}{2}\,\omega^{-1}\biggl{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\biggr{)}.$
-Furthermore, for each $x\in{\mathcal{K}}$ and $i=1,\dots,L$, if $\|x-v_{i}\|<\delta$ then
+Let ${\mathcal{K}}$ be a compact subset of $\mathbb{R}^{n}$ whose doubling number is $C$.
+Fix an “number of experts-to-expert complexity trade-off parameter” $r\in\mathbb{R}$ and a “valency parameter” $\nu\in\mathbb{N}$ with $\nu\geq 2$.
+For every uniformly continuous map $f:{\mathcal{K}}\to\mathbb{R}^{m}$ with modulus of continuity $\omega$ and each approximation error $\varepsilon>0$, $p\in\mathbb{N}_{+}$, there is an $\nu$-ary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq{\mathcal{K}}\times\mathbb{R}^{p}$ and a family of MLPs with (P)ReLU activation function $\{\hat{f}_{\theta_{i}}\}_{i=1}^{L}$ defined by $p$ parameters mapping $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$ satisfying:
+$\max_{x\in{\mathcal{K}}}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|<\frac{\varepsilon^{-2r/n}}{2}\,\omega^{-1}\biggl{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\biggr{)}$
+and for each $x\in{\mathcal{K}}$ and $i=1,\dots,L$, if $\|x-v_{i}\|<\delta$ then
 $\|f(x)-f_{\theta_{i}}(x)\|<\varepsilon.$
 The depth and width of each $\hat{f}_{\theta_{i}}$ and the number of leaves, height, and number of nodes required to build the $\nu$-ary tree are all recorded in Table 1.
-Table 6: Complexity of Feedforward Neural Network $\hat{f}_{\theta}$ and the $\nu$-ary tree in Theorem A.5. Here $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
+Table 4: Complexity of Feedforward Neural Network $\hat{f}_{\theta}$ and the $\nu$-ary routing tree in Theorem 5.3. Here $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
+Parameter Estimate
+Depth $(J)$ $\mathcal{O}(\max\{1,\varepsilon^{-r}\})$
+Width ($\max_{j}\,d_{j}$) $\mathcal{O}(1)$
+No. Experts $\mathcal{O}\big{(}\max\{1,\varepsilon^{2r/n}/\omega^{-1}(\varepsilon)\}\big{)}$
+Routing Complexity $\mathcal{O}\big{(}\max\{1,\log(\varepsilon^{2r/n}/\omega^{-1}(\varepsilon))\}\big{)}$
+See Table 6 in Appendix A for more detailed estimates.
+6 Conclusion
+We presented approximation-theoretic and statistical foundations for MoEs by analysing the MoMLP model.
+We found that MoMLPs can indeed achieve arbitrary uniform approximation accuracy of continuous functions on compact subsets of Euclidean space while maintaining a feasible number of parameters in active GPU VRAM memory (Theorem 5.3).
+However, this naturally comes at the cost of requiring an exponential number of experts.
+We also obtain upper bounds on the VC dimension of the MoMLP model (Theorem 4.2), akin to the results of Bartlett et al.
+(2019) for ReLU MLPs, showing that deep-learning-based MoEs can indeed generalize.
+Limitations and Future Work
+This paper provides theoretical support, both in terms of approximation theory (approximation rates) and learning theory (VC dimension bounds), of MoE models.
+As such there is no clear limitation to our analysis.
+In future work, one can use the tools developed here to similarly study MoEs of other (deep) learning models.
+Societal Impact
+This paper presents work aimed at advancing the theoretical foundations of machine learning.
+Therefore, the majority of societal consequences are more positive than negative, as they acknowledge the necessity for principled machine learning and contribute to the development of understanding the capabilities of AI.
+Acknowledgment and Funding
+The authors thank James Lucas, Rafid Mahmood, and Gabriel Conant for their helpful feedback while preparing the manuscript.
+AK acknowledges financial support from an NSERC Discovery Grant No.
+RGPIN-2023-04482 and their McMaster Startup Funds.
+AK also acknowledges that resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute https://vectorinstitute.ai/partnerships/current-partners/.
+HSOB acknowledges financial support from an EV travel grant.
+Appendix A Detailed Tables And Rates
+Table 5: Complexity of the MLP $\hat{f}_{\theta}$ in Lemma 5.2.
+Activation $\sigma$ Super Expressive 4.3 PReLU 3.1
+[…]
+Width ($\max_{j}\,d_{j}$) $36n(2n+1)+m$ $16\max\{n,3\}+m$
+Table 6: Complexity of Feedforward Neural Network $\hat{f}_{\theta}$ and the $\nu$-ary tree in Theorem 5.3. Here $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
 Parameter Estimate
@@
 Width ($\max_{j}\,d_{j}$) $16\max\{n,3\}+m$
-Leaves $\mathcal{O}\Big{(}v^{\big{\lceil}c\,\log(C)\big{(}1+\log(\varepsilon^{2r/n}\,\operatorname{diam}({\mathcal{K}}))/\big{(}2\,\omega^{-1}\big{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\big{)}\big{)}\big{)}\big{\rceil},}\Big{)}$
-Height $\big{\lceil}c\,\log(C)\big{(}1+\log(\epsilon^{2r/n}\operatorname{diam}({\mathcal{K}})/\big{(}2\omega^{-1}\big{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\big{)}\big{)})\big{)}\big{\rceil}$
+No.
+Experts (No.
+Leaves) $\mathcal{O}\Big{(}v^{\big{\lceil}c\,\log(C)\big{(}1+\log(\varepsilon^{2r/n}\,\operatorname{diam}({\mathcal{K}}))/\big{(}2\,\omega^{-1}\big{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\big{)}\big{)}\big{)}\big{\rceil},}\Big{)}$
+Height (Routing Complexity) $\big{\lceil}c\,\log(C)\big{(}1+\log(\epsilon^{2r/n}\operatorname{diam}({\mathcal{K}})/\big{(}2\omega^{-1}\big{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\big{)}\big{)})\big{)}\big{\rceil}$
 Nodes $\mathcal{O}\Biggl{(}\frac{v^{\lceil c\,\log(C)(1+\log(\varepsilon^{2r/n}\operatorname{diam}({\mathcal{K}})/\big{(}2\omega^{-1}(\frac{\varepsilon}{131\,(nm)^{1/2}})\big{)}))\rceil+1}-1}{\big{\lceil}c\log(C)\big{(}1+\log(\frac{\varepsilon^{2r/n}}{2}\operatorname{diam}({\mathcal{K}}))/\omega^{-1}\biggl{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\biggr{)}\big{)}\big{\rceil}-1}\Biggr{)}$
-Appendix B Appendix: Proof of Main Theorem
-Proof of Lemma A.2.
+Appendix B Appendix: Proofs
+B.1 Proof of Theorem 4.1
+Proof of Lemma 5.1.
 Since ${\mathcal{K}}$ is a doubling metric space then, (Acciaio et al., 2023, Lemma 7.1), for each $\delta>0$, there exist $x_{1},\dots,x_{N}\in\mathcal{K}$ satisfying
@@
 The original formulation of that result only allows for $\alpha$-Hölder function.
-Proof of Lemma A.3.
+Proof of Lemma 5.2.
 If $f(x)=c$ for some constant $c>0$, then the statement holds with the neural network $\hat{f}(x)=c$, which can be represented as in (2) with $[d]=(n,m)$, where $A^{j}$ is the 00 matrix for all $j$, and the “$c$” in (2) is taken to be this constant $c$.
@@
 $r_{\mathcal{K}}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{diam}({\mathcal{K}})\sqrt{\frac{n}{2(n+1)}}.$ (14)
-By Jung’s Theorem (see (Jung, 1901)), there exists $x_{0}\in\mathbb{R}^{n}$ such that the closed Euclidean ball $\overline{\operatorname{Ball}_{(\mathbb{R}^{n},d_{n})}\left(x_{0},r_{\mathcal{K}}\right)}$ contains ${\mathcal{K}}$.
+By Jung’s Theorem (see Jung (1901)), there exists $x_{0}\in\mathbb{R}^{n}$ such that the closed Euclidean ball $\overline{\operatorname{Ball}_{(\mathbb{R}^{n},d_{n})}\left(x_{0},r_{\mathcal{K}}\right)}$ contains ${\mathcal{K}}$.
 Therefore, by Hölder’s inequality, we have that the $n$-dimensional hypercube $[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]$   [^7] For $x,y\in\mathbb{R}^{n}$ we denote by $[x,y]$ the hypercube defined by $\prod_{i=1}^{n}[x_{i},y_{i}]$.contains $\overline{B_{({\mathbb{R}^{n}},d_{n})}\left(x_{0},r_{\mathcal{K}}\right)}$, where $\bar{1}=(1,\dots,1)\in\mathbb{R}^{n}$.
@@
   1. (i)
-    If $\sigma$ satisfies Definition A.1 then, setting each $\alpha=0$ implies that $\sigma_{0}(x)=\sigma^{\star}(x)$, as defined in Eq. (10); thus
+    If $\sigma$ satisfies Definition 4.3 then, setting each $\alpha=0$ implies that $\sigma_{0}(x)=\sigma^{\star}(x)$, as defined in Eq. (8); thus
     $J^{(i)}\leq 11\,\mbox{ and }\,\max_{1\leq j\leq J^{(i)}}d_{j}\leq 36n(2n+1)$
@@
 Let $g_{1}\bullet g_{2}$ denote the component-wise composition of a univariate function $g_{1}$ with a multivariate function $g_{2}$.
-If the activation function $\sigma$ is either in Definitions A.1 or 3.1, then it trivially implements the identity $I_{\mathbb{R}}$ on $\mathbb{R}$ by setting $\alpha=0$; i.e.
+If the activation function $\sigma$ is either in Definitions 4.3 or 3.1, then it trivially implements the identity $I_{\mathbb{R}}$ on $\mathbb{R}$ by setting $\alpha=0$; i.e.
 $\sigma_{1}=I_{\mathbb{R}}$.
@@
   1. (i)
-    If $\sigma$ satisfies Definition A.1 then, setting each $\alpha=0$
+    If $\sigma$ satisfies Definition 4.3 then, setting each $\alpha=0$
     $J\leq$ $15\,m$
@@
 Incorporating the definition of $r_{\mathcal{K}}$ in Eq. (14) and employing the inequality $\operatorname{diam}({\mathcal{K}})\leq 2\operatorname{rad}({\mathcal{K}})$ completes the proof. ∎
-Proof of Lemma A.4.
-Consider the tree $\tilde{\mathcal{T}}$ given by Lemma A.2.
-For each leaf $v_{i}$ of $\tilde{\mathcal{T}}$, we apply Lemma A.3 to deduce the existence of an MLP $\hat{f}_{\theta_{i}}$ with explicit depth and width estimates given by that lemma, satisfying the uniform estimate
+Lemma B.1 (Trade-Off: No. Expert vs. Expert Complexity).
+Let ${\mathcal{K}}$ be a compact subset of $\mathbb{R}^{n}$ whose doubling number is $C$ and a uniformly continuous map $f:{\mathcal{K}}\to\mathbb{R}^{m}$ with modulus of continuity $\omega$.
+Fix $v\in\mathbb{N}$ with $v\geq 2$, $0<\delta\leq\operatorname{rad}({\mathcal{K}})$, and $\varepsilon>0$.
+Suppose that $\sigma$ satisfies Definition 3.1.
+There exists a $p\in\mathbb{N}_{+}$ and a $v$-ary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq{\mathcal{K}}\times\mathbb{R}^{p}$ satisfying
+$\max_{x\in{\mathcal{K}}}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|<\delta.$ (18)
+Furthermore, for each $x\in{\mathcal{K}}$ and each $i=1,\dots,L$, if $\|x-v_{i}\|<\delta$ then
+$\|f(x)-f_{\theta_{i}}(x)\|<\varepsilon.$
+We have the following estimates:
+  1. (i)
+    Depth.
+Depth of each $\hat{f}_{\theta_{i}}$ is $m\left(19+2n+11\Big{\lceil}\Big{(}\frac{\delta 2^{3/2}n^{1/2}}{(n+1)^{1/2}\omega^{-1}\big{(}\varepsilon/(131\sqrt{n\,m})\big{)}}\Big{)}^{n/2}\Big{\rceil}\right)$
+  2. (ii)
+    Width. Width of each $\hat{f}_{\theta_{i}}$ is $16\,\max\{n,3\}+m$
+  3. (iii)
+    Leaves: at most $L=v^{\big{\lceil}c\,\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil},}$
+  4. (iv)
+    Height: $\big{\lceil}c\,\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil}$,
+  5. (v)
+    Nodes: At most $\frac{v^{\lceil c\,\log(C)(1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}})))\rceil+1}-1}{\big{\lceil}c\log(C)\big{(}1+\log(\delta^{-1}\operatorname{diam}({\mathcal{K}}))\big{)}\big{\rceil}-1}$
+where $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
+Proof of Lemma B.1.
+Consider the tree $\tilde{\mathcal{T}}$ given by Lemma 5.1.
+For each leaf $v_{i}$ of $\tilde{\mathcal{T}}$, we apply Lemma 5.2 to deduce the existence of an MLP $\hat{f}_{\theta_{i}}$ with explicit depth and width estimates given by that lemma, satisfying the uniform estimate
 $\max_{\|x-v_{i}\|\leq\delta}\,\|f(x)-\hat{f}_{\theta_{i}}(x)\|<\varepsilon.$ (19)
 Let $\mathcal{T}$ be the same tree as $\tilde{\mathcal{T}}$ with leaves identified with $\{(v_{i},\theta_{i})\}_{i=1}^{L}$. ∎
-We now prove our main technical version, namely Theorem A.5, which directly implies the special case recorded in Theorem 3.4.
-Proof of Theorem A.5 (and this Theorem 3.4).
-Applying Lemma A.4 with $\delta$ given as the solution of
+We now prove our main technical version, namely Theorem 5.3, which directly implies the special case recorded in Theorem 4.1.
+Proof of Theorem 5.3 (and this Theorem 4.1).
+Applying Lemma B.1 with $\delta$ given as the solution of
 $\Big{(}\frac{\delta 2^{3/2}n^{1/2}}{(n+1)^{1/2}\omega^{-1}(\varepsilon/131\,(nm)^{1/2})}\Big{)}^{n/2}\leq\Big{(}\frac{\delta 2^{3/2}n^{1/2}}{(2n^{1/2}\omega^{-1}(\varepsilon/131\,(nm)^{1/2})}\Big{)}^{n/2}=\varepsilon^{-r}.$ (20)
@@
 Proof of Theorem 4.1.
-Setting ${\mathcal{K}}=B_{n}(0,1)$, $r=1/2$, $\omega(t)=Lt$, and thus $\omega^{-1}(t)=L^{-1}\,t^{1/\alpha}$, in Theorem A.5 yields the conclusion.
+Setting ${\mathcal{K}}=B_{n}(0,1)$, $r=1/2$, $\omega(t)=Lt$, and thus $\omega^{-1}(t)=L^{-1}\,t^{1/\alpha}$, in Theorem 5.3 yields the conclusion.
 Finally, by the computation in Example 1, we have that $C\leq 2^{n+1}$; thus, $\log(C)=(n+1)\log(2)\leq 2n$.
@@
 ∎
-Remark B.1.
-[…]
-B.1 Proofs of Results Concerning VC dimension
+B.2 Proofs of Theorem 4.2
 We use the following lemma and its proof due to Gabriel Conant.
-Lemma B.2 ((Conant, )).
+Lemma B.2 (Conant (2023)).
 Fix $n,L,d\in\mathbb{N}_{+}$.
@@
 ∎
-We may now derive Proposition 3.6 by merging Lemma B.2 and one of the main results of Bartlett et al. (2019).
-Proof of Proposition 3.6.
+We may now derive Theorem 4.2 by merging Lemma B.2 and one of the main results of Bartlett et al. (2019).
+Proof of Theorem 4.2.
 Let $n,J,W,L\in\mathbb{N}_{+}$ and consider the (non-empty) set of real-valued functions $\mathcal{NN}_{J,W:n,1}^{\operatorname{PReLU}}$.
@@
 yielding the second conclusion. ∎
-Proof of Proposition 3.7.
+B.3 Proof of Proposition 4.4
+Proof.
 We argue by contradiction.
@@
 (4), we infer that the minimum number layers $(L)$ and minimal width $(W)$ of each MLP approximating a low-regularity function on the low-dimensional domain $[0,1]$ is roughly the same as the minimal number of layers and width of an MLP approximating a high-regularity map on the high-dimensional domain $[0,1]^{n}$.
-Appendix D Experimental details
+Appendix D Does one Need Overparameterized Experts if there are Enough Experts?
+We evaluate our approach in two standard machine learning tasks: regression and classification.
+We experimentally show that MoMLPs, which distribute predictions over multiple neural networks, are competitive with a single large neural network containing as many model parameters as all the MoMLPs combined.
+This is desirable in cases where the large neural network does not fit into the memory of a single machine.
+On the contrary, the MoMLP model can be trained by distributing each MoMLP on separate machines (or equivalently serially on a single machine).
+Inference can then be performed by loading only a single MoMLP at a time into the GPU.
+Our source code to reproduce our results is available in the supplementary material.
+D.1 Regression
+We first consider regression, where the goal is to approximate non-convex synthetic functions often used for performance test problems.
+In particular, we choose 1-dimensional Hölder functions, as well as Ackley Ackley (1987) and Rastrigin Rastrigin (1974) functions, whose formulations are detailed in Appendix D.3.
+1D Hölder Functions.
+We illustrate our primary finding, as encapsulated in Theorem 4.1, by leveraging $1$-dimensional functions characterized by very low regularity.
+This choice is motivated by the jagged structure inherent in such functions, necessitating an exponentially higher sampling frequency compared to smooth functions for achieving an accurate reconstruction.
+Indeed, this crucial sampling step forms the foundation of many quantitative universal approximation theorems Yarotsky (2017); Shen et al.
+(2021); Kratsios & Papon (2022).
+As elaborated in Appendix C, approximating a well-behaved (Lipschitz) function in $d$ dimensions poses a challenge equivalent to approximating a highly irregular function ($1/d$-Hölder) in a single dimension.
+A visual representation of a $1/d$-Hölder function is presented in Figure 2, exemplified by the trajectory of a fractional Brownian motion with a Hurst parameter of $\alpha=1/d$.
+A formal definition is available in Appendix C.
+Figure 2: Visual Comparison of Functions with High ($\alpha\approx 1$) vs.
+Low ($\alpha\approx 0$) Hölder regularity.
+If $\alpha\approx 1$, the function (green) is approximately differentiable almost everywhere, meaning it does not osculate much locally and thus is simple to approximate.
+If $\alpha\approx 0$, the function may be nowhere differentiable and jagged; its extreme details make it difficult to approximate.
+Table 7: Test mean squared error (average and standard deviation) for the different functions of the regression task.
+1D Hölder 2D Ackley 3D Ackley 2D Rastrigin 3D Rastrigin
+MoMLPs (ours) 0.057 $\pm$ 0.085 0.00015 $\pm$ 0.00006 0.00068 $\pm$ 0.00010 0.0480 $\pm$ 0.0073 1.0062 $\pm$ 0.0446
+Baseline 0.128 $\pm$ 0.012 0.08723 $\pm$ 0.01059 0.09303 $\pm$ 0.03156 3.0511 $\pm$ 0.3581 8.0376 $\pm$ 4.0499
+2D and 3D Functions. We select the Ackley and Rastrigin functions, with their respective 2D representations showcased in Figure 3, as widely recognized benchmarks in the field of optimization.
+(a) Ground Truth Ackley
+(b) Ground Truth Rastrigin
+(c) Predicted Ackley
+(d) Predicted Rastrigin
+(e) log MSE
+(f) log MSE
+Figure 3: Comparison of ground truth and predicted results for 2D Ackley and Rastrigin functions over the domain $[-1,1]^{2}$.
+Evaluation protocol.
+We consider the setup where the domain of a function that we try to approximate is the $n$-dimensional closed set $[a,b]^{n}$.
+For instance, we arbitrarily choose the domain $[0,1]$ when $n=1$, and $[-1,1]^{n}$ when $n\geq 2$.
+Our training and test samples are the $s^{n}$ vertices of the regular grid defined on $[a,b]^{n}$.
+At each run, $80\%$ of the samples are randomly selected for training and validation, and the remaining $20\%$ for testing.
+During training, for a given and fixed set of $K$ prototypes $p\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(p_{1},\dots,p_{K})$, we assign each training sample $x$ to its nearest prototype $p_{k}$ and associated neural network $\hat{f}_{k}$.
+We learn the prototypes as explained in Section D.4.
+For simplicity, we set the number of prototypes to $K=4$; we also set $s=10,000$ if $n=1$, $s=150$ if $n=2$ (i.e., $500150^{2}=22,$ 500 samples), and $s=30$ if $n=3$ (i.e., $27,000$ samples).
+More details can be found in Appendix D.3.
+Test performance.
+In Table 7, we present the mean squared error obtained on the test set across 10 random initializations and various splits of the training/test sets.
+The baseline consists of a single neural network with the same overall architecture as each MoMLP but possesses as many parameters as all the MoMLPs combined.
+In all instances, the MoMLP model demonstrates a significant performance advantage compared to the baseline.
+Figure 3 illustrates the predictions generated by our MoMLPs, showcasing the capability of our approach to achieve a good approximation of the ground truth functions.
+D.2 Classification
+Datasets.
+We evaluate classification on standard image datasets such as CIFAR-10 (Krizhevsky & Hinton, 2010), CIFAR-100, and Food-101 (Bossard et al., 2014), which consist of 10, 100, and 101 different classes, respectively.
+We use the standard splits of training/test sets: the datasets include (per category) 5,000 training and 1,000 test images for CIFAR-10, 500 training and 100 test images for CIFAR-100, and 750 and 250 for Food-101.
+Training.
+Our MoMLP model takes as input latent DINOv2 encodings (Oquab et al., 2023) of images from the aforementioned datasets.
+Each sample $x\in\mathbb{R}^{768}$ corresponds to a DINOv2 embedding (i.e., $n=768$).
+Additionally, we set the prototypes as centroids obtained through the standard $K$-means clustering on the DINOv2 embedding space.
+The replacement of our original prototype learning algorithm is sensible in this context, as we operate within a structured latent space optimized through self-supervised learning using large-scale compute and datasets.
+Table 8: Test classification accuracy using DINOv2 features as input (average and standard deviation).
+Dataset CIFAR-10 CIFAR-100 Food-101
+Ours (weighted) 98.40 $\pm$ 0.05 90.01 $\pm$ 0.11 91.86 $\pm$ 0.10
+Ours (unweighted) 98.42 $\pm$ 0.04 89.62 $\pm$ 0.25 91.79 $\pm$ 0.16
+Baseline 98.45 $\pm$ 0.06 89.85 $\pm$ 0.17 91.45 $\pm$ 1.09
+Due to the potential class imbalance in the various Voronoi cells formed by the prototypes, we utilize two variations of the cross-entropy loss for each MoMLP $\hat{f}_{k}$.
+The first variation, termed unweighted, assigns equal weight to all categories.
+The weighted variation assigns a weight that is inversely proportional to the distribution of each category in the Voronoi cell defined by the prototype $p_{k}$.
+Test performance.
+We present the test classification accuracy over 10 random initializations of both the baseline and our MoMLPs in Table 8.
+The weighted version performs slightly better on datasets with a large number of categories.
+Our approach achieves comparable results with the baseline, effectively decomposing the prediction across multiple smaller models.
+D.3 Experimental details
 We include here experimental details, we refer to the source code in the supplementary material for more details.
-D.1 Formulations of the Ackley and Rastrigin functions
+We first outline the algorithm used to train the MoMLP MoE model.
+We then provide details on the trained architecture and hyperparameter details in the implementation.
+D.4 Training Algorithm
+We now provide an explanation for the training algorithm.
+As discussed, we mitigate down the algorithm into two parts: discovering the prototypes and training the networks.
+Conceptually, the prototypes define where in the input space the networks are located, or in other words, where in the input space we expect each of the networks to have the best performance.
+During inference, we will route a given input to the appropriate network based on its nearest prototype.
+In essence, each network learns to approximate a specific region of the overall input domain.
+Discovering prototypes.
+In principle, we may not know how to partition the input space.
+One approach is to utilize standard clustering algorithms like $K$-means, but this might be suboptimal for the downstream task unless we are already operating in a structured latent space, such as those found in pre-trained models (further discussion on this is available in Section D).
+Another way is to learn it via gradient descent by optimizing the location of the prototypes for a specific task by following the gradient of the downstream loss.
+At the beginning of training, we have $\hat{F}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\hat{f}_{1}(x),\dots,\hat{f}_{K}(x))$ which contains a collection of $K$ randomly initialized shallow or small networks (i.e., much smaller than our MoMLPs described later).
+In this first step, we assume that we are able to load all randomly initialized networks into our GPU memory.
+In particular, this is true because we use small networks with few parameters, which we will later “deepen” in the next step by adding additional hidden layers.
+We initialize the prototypes $p\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(p_{1},\dots,p_{K})$ randomly from a uniform distribution within the bounds of our training dataset input samples.
+We use the following expression to train the location of our prototypes $\{p_{k}\}_{k=1}^{K}$ in $\mathbb{R}^{n}$ by minimizing the energy:
+$\sum_{(x,y)\in\mathcal{D}}\,\ell\big{(}\operatorname{softmax}\big{(}-\|x-p_{i}\|_{i=1}^{K}\big{)}^{\top}\,\hat{F}(x)\,,y\big{)}.$ (36)
+where the loss $\ell$ is task-specific; for example, one could use mean squared error for regression and cross-entropy for classification.
+The $\operatorname{softmax}$ weights the importance of the prediction of each of the MoMLPs in $\hat{F}$ for a given input $x$, as a function of the input’s distance to the prototypes, $\|x-p_{i}\|_{i=1}^{K}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\|x-p_{1}\|,...,\|x-p_{K}\|)$.
+Both the locations of the prototypes and the shallow randomly initialized neural networks assigned to them are optimized.
+Deepening the Networks.
+After the initial training phase, we enhance the networks by incorporating additional layers.
+Specifically, we introduce linear layers with weights initialized to the identity matrix and bias set to zero, just before the final output layer of each network.
+To encourage gradient flow in these new layers during the subsequent training stage, we slightly perturb this initialization with small Gaussian noise.
+This approach is driven by the fact that in the second training stage, each MoMLP can be optimized in a distributed manner.
+Consequently, we can work with larger networks without the need to load all of them simultaneously into our GPU, allowing for more model parameters.
+During the first stage, we have already optimized our networks alongside the prototype locations, converging towards a minimum.
+By initializing the networks with the new layers close to the identity, we can ensure that their output at the start of the second stage of training is similar to that produced by the original networks.
+This allows us to smoothly continue the optimization process from the point where we previously halted.
+MoMLP Training. Once prototype locations have been fixed we can independently train MoMLPs $\hat{f}_{1},\dots,\hat{f}_{K}$ by minimizing for all $k\in\{1,\dots,K\}$:
+($\underset{\underset{k\in\operatorname*{arg\,min}_{j\in\{1,\dots,K\}}\{\|x-p_{j}\|\}}{(x,y)\in\mathcal{D}}}{\sum}\,\ell(\hat{f}_{k}(x),y)$ (37)
+over all the networks. We optimize the MoMLP network $\hat{f}_{k}$ for training data points that are closest to prototype $p_{k}$. The training procedure is summarized in Algorithm 1.
+1Training data $\mathcal{D}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(x_{j},y_{j})\}_{j=1}^{N}$, no. of prototypes $K\in\mathbb{N}_{+}$, loss function $\ell$.
+Discovering Prototypes:
+2
+3
+$(\hat{F},p)\leftarrow{\operatorname*{arg\,min}_{\hat{F},p}}\,\displaystyle\sum_{(x,y)\in\mathcal{D}}\,\ell\left(\operatorname{softmax}\left(x|p\right)^{\top}\hat{F}(x),y\right)$
+Deepen networks:
+    For $k=1,\dots,K$:
+4     $\hat{f}_{k}\leftarrow\textrm{deepen}(\hat{f}_{k})$
+MoMLP Training:
+   For $k=1,\dots,K$:
+5     $\hat{f}_{k}\leftarrow\underset{\hat{f}_{k}}{\operatorname*{arg\,min}}\,\underset{\underset{k\in\operatorname*{arg\,min}_{j}\{\|x-p_{j}\|\}}{(x,y)\in\mathcal{D}}}{\sum}\,\ell(\hat{f}_{k}(x),y)$
+return MoMLP parameters $\{\hat{f}_{k}\}_{k=1}^{K}$ and prototype locations $\{p_{k}\}_{k=1}^{K}$.
+Algorithm 1 MoMLPs Training.
+Inference.
+At inference time, each test sample $x$ is assigned to its nearest prototype $p_{k}$ where $k\in\operatorname*{arg\,min}_{j\in\{1,\dots,K\}}\{\|x-p_{j}\|\}$ and the prediction is made by the $k$-th MoMLP $\hat{f}_{k}$.
+Comparison to Standard Distributed Training.
+One can distribute the complexity of feedforward models by storing each of their layers in offline memory and then loading them sequentially into VRAM during the forward pass.
+This does avoid loading more than $\mathcal{O}(\operatorname{Width})$ active parameters into VRAM at any given time, where $\operatorname{Width}$ denotes the width of the feedforward model.
+However, doing so implies that all the model parameters are ultimately loaded during the forward pass.
+This contrasts with the MoMLP model, which requires $\mathcal{O}\big{(}\log_{2}(K)\,\operatorname{Width}^{2}\,\operatorname{Depth}\big{)}$ to be loaded into memory during a forward pass; where $\operatorname{Width}$ and $\operatorname{Depth}$ are respectively the largest width and depth of the MLP at any leaf of the tree defining a given MoMLP, and $K$ denotes the number of prototypes.
+However, in the forward pass, one loads $\mathcal{O}(\varepsilon^{-n/2})$ parameters for the best worst-case MLP while only $\mathcal{O}(n\log(1/\varepsilon)/\varepsilon)$ are needed in the case of the MoMLPs.
+The number of parameters here represents the optimal worst-case rates for both models (see Table 6 and Theorem 5.3).
+D.5 Definitions of the Ackley and Rastrigin functions
 Let us note $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$ the $n$-dimensional representation of a sample, we use the following formulation of the Ackley function:
@@
 $\operatorname{Rastrigin}(x)=\sum_{i=1}^{n}x_{i}^{2}+10\left(n-\sum_{i=1}^{n}\cos(2\pi x_{i})\right)$ (39)
-D.2 Architectures and hyperparameters
-D.2.1 Neural pathways
-We consider that the width of our neural pathways is $w=1000$. In other words, each hidden layer of our neural pathways contains a linear matrix of size $w\times w$.
-In the regression task, our neural pathways contain 3 hidden layers and we use a learnable PReLU as activation function.
-We use Adam (Kingma & Ba, 2014) as optimizer with a learning rate of $410^{-}$ 4 and other default parameters of Pytorch.
+D.6 Architectures and hyperparameters
+D.6.1 MoMLPs
+We consider that the width of our MoMLPs is $w=1000$. In other words, each hidden layer of our MoMLPs contains a linear matrix of size $w\times w$.
+In the regression task, our MoMLPs contain 3 hidden layers and we use a learnable PReLU as activation function.
+We use Adam Kingma & Ba (2014) as optimizer with a learning rate of $410^{-}$ 4 and other default parameters of Pytorch.
 In the classification task, we follow the setup of Oquab et al.
-(2023) and then use AdamW (Loshchilov & Hutter, 2019) as optimizer with a learning rate of $10^{-3}$ and other default parameters of Pytorch, our neural pathways contain 4 hidden layers in the classification task, and we apply BatchNorm1d before the PReLU activation function.
-D.2.2 Baseline
-The baseline has the same architecture as the neural pathways above.
-However, if we note $K$ the number of prototypes and assume that $\sqrt{K}$ is a natural number, the width of the baseline is $w\sqrt{K}$ so that the number of hidden parameters is the same as all the neural pathways combined.
+(2023) and then use AdamW Loshchilov & Hutter (2019) as optimizer with a learning rate of $10^{-3}$ and other default parameters of Pytorch, our MoMLPs contain 4 hidden layers in the classification task, and we apply BatchNorm1d before the PReLU activation function.
+D.6.2 Baseline
+The baseline has the same architecture as the MoMLPs above.
+However, if we note $K$ the number of prototypes and assume that $\sqrt{K}$ is a natural number, the width of the baseline is $w\sqrt{K}$ so that the number of hidden parameters is the same as all the MoMLPs combined.
 In our experiments, we set $K=4$, so $\sqrt{K}=2$.